### v2

# ConceptDrift: Uncovering Biases through the Lens of Foundation Models

Abstract
An important goal of ML research is to identify and mitigate unwanted biases intrinsic to datasets and already incorporated into pre-trained models. Previous approaches have identified biases using highly curated validation subsets, that require human knowledge to create in the first place. This limits the ability to automate the discovery of unknown biases in new datasets. We solve this by using interpretable vision-language models, combined with a filtration method using LLMs and known concept hierarchies. More exactly, for a dataset, we use pre-trained CLIP models that have an associated embedding for each class and see how it drifts through learning towards embeddings that disclose hidden biases. We call this approach ConceptDrift and show that it can be scaled to automatically identify biases in datasets like ImageNet without human prior knowledge. We propose two bias identification evaluation protocols to fill the gap in the previous work and show that our method significantly improves over SoTA methods, both using our protocol and classical evaluations. Alongside validating the identified biases, we also show that they can be leveraged to improve the performance of different methods. Our method is not bounded to a single modality, and we empirically validate it both on image (Waterbirds, CelebA, ImageNet), and text datasets (CivilComments).
1 Introduction
Deep neural networks, and especially fine-tuned versions of foundation models, are commonly deployed in critical areas such as healthcare, finance, and criminal justice, where biased predictions can have significant societal consequences [1]. Despite their impact, these models are often employed in their natural black-box state, i.e. as highly non-linear, multi-layered decision processes, lacking transparency or interpretability. Even if the pretrained model has been validated by the community, the dataset leveraged in the fine-tuning process can, and usually does, imprint the model with new biases. This issue is particularly concerning as biases can lead to undesired outcomes [6], reinforcing existing inequalities or creating new forms of discrimination.
Subpopulation shift setups [40, 21, 5], featuring naturally occurring spurious correlations, provided researchers with a controlled environment for studying bias detection and prevention. Within this context, recent efforts began employing foundation models [15, 47] in their investigation protocols. However, current methods limit themselves to data analysis alone. For instance, the method proposed by Kim et al. [15] relies on validating the presence of a given object within the set of mistakes and its absence from within the set of the correctly classified samples within the validation set, in order to label it as a bias. Thus, data-restricted methods rely on the critical assumption that the validation set is exhaustive in some sense, that it contains samples capable of exposing all biases. This assumption is justified only within the context of a controlled experimental environment, such as having a subpopulation shift setup where all sources of bias are given or known. Any other real-world scenario hardly guarantees an exhaustive validation set.
Figure 1: Following ConceptDrift’s steps for the ImageNet ”Fire Truck” class. In Step 1, during training, the classification weights $W$ drift from the initial class concept embedding $W^{0}$, outside the scope of relevant concepts, towards biases. In Step 2, our method filters out class-related concepts and, using an embedding-space scoring system, ranks and automatically selects the highest-ranking class-neutral concepts as biases. Leveraging these identified biases, we generate synthetic images that are able to fool even strong pre-trained methods.
We expand the current usage of foundation models beyond the restricted scope of data analysis and propose a new direction for bias identification. Our method focuses on investigating the skewness of the model’s weights towards detecting and prioritizing spurious features as part of the decision-making process of the investigated model. We hereby propose a novel protocol for uncovering biases using foundation models such as CLIP [30] and mGTE [45], leveraging the topology of their embedding space to identify and name biases instilled by linear probing. Our protocol, dubbed ConceptDrift, is illustrated in Fig. 1. We showcase how, during training, the weights of the final classification layer drift away from the textual representation of their associated class, towards representations of spurious attributes. We propose a ranking system based on embedding-space arithmetic to extract concepts that factor in the activation of class neurons and delineate the highest-ranking concepts that lie outside the semantic scope of the classes, as biases.
Our main contributions are summarized as follows:
  1. 1.
    We introduce ConceptDrift, a method capable of pin-pointing concepts that are relevant for the decision-making process of a model. We are the first to propose a weight-space approach for identifying biases that a foundation model learns from a given dataset, diverging from the current techniques oriented on validation set errors.
  2. 2.
    We propose two novel evaluation strategies for the bias identification task. The first one evaluates the performance decay over generated samples that exploit the identified biases, highlighting the potential harm biases can cause to models. The second one assesses the performance gains achieved through bias-based regularization in a training setup with a 100% spurious correlation between classes and biases, focusing on mitigating biases and enhancing model robustness.
  3. 3.
    We validate ConceptDrift, using both established evaluation protocols and our newly introduced strategies. We achieve significant advances upon the SoTA bias identification techniques on classical image datasets: Waterbirds [40] and CelebA [21], and further extend the evaluation by including a text dataset: CivilComments [5]. We also quantitatively confirm biases identified in ImageNet with respect to a foundation model.
  4. 4.
    We show that more robust methods can be obtained by devising constraints around the discovered biases. This includes subgroup constraints, training regularization, and zero-shot prompting.
Table 1: Comparison for Bias Identification approaches, based on fundamental differences in methods.
Method Root of bias identification Scoring bias candidates Fully auto Evaluation on large datasets Evaluation protocols employed
SpLiCE [4] disentangled representations Lasso solver Masking concepts from decomposition
B2T [15] mistakes in valid-set keywords frequency GroupDRO over bias-based groups
Lg [47] class-specificity score latent-space arithmetic GroupDRO over bias-based groups Train on bias-exploiting samples
ConceptDrift (ours) biases’ similarity to class weights latent-space arithmetic partially Validate on bias-exploiting samples (Sec. 5.1) GroupDRO over bias-based groups (Sec. 5.2) Bias-based training regularization (Sec. 5.3)
2 Related Work
Bias identification for generalization Machine learning methods naturally capture relevant factors needed to solve a task. Nevertheless, many times, models also capture shortcuts [12], that are helpful in solving the task, but are not fundamental or essential for it. These shortcuts represent spurious correlations or biases, that don’t always hold, and should not be used for reliable generalization outside the training distribution, often leading to degraded performance [29, 3, 13]. Prior works [15, 47, 4] have thus focused on identifying dataset biases, mainly through data analysis procedures. Kim et al. [15] identify concepts that are more prevalent in the misclassified examples in comparison with the correctly classified ones, while Zhao et al. [47] look for concepts that are specific to one class. In contrast, Bhalla et al. [4] identifies biases by training a linear layer on top of a disentangled representation of the data and having humans decide which features should not have been used for the task. For a meaningful comparison, we show in Tab. 1 existing bias identification methods, highlighting aspects we consider fundamental for the bias identification task.
Debiasing Debiasing and bias extraction techniques have become crucial in ensuring the fairness and accuracy of machine learning models [36], with extensive research dedicated to removing harmful biases across various domains. Some existing methods use bias annotations to train unbiased models, by means of group balanced subsampling [14], reweighting [33] or data augmentations [44]. In the absence of these annotations, other works [25, 20, 28] have proposed to first learn a biased model and then focus on its mistakes to train an unbiased one.
Fairness Fairness in has been extensively studied, with numerous approaches [8, 38] proposed to facilitate ethical research and to ensure equitable outcomes across different subpopulations. Most of those methods overlap with domain generalization and worst-group performance improvements. This is also a field where model interpretability plays a crucial role [32], as understanding how decisions are made can help in identifying and mitigating biases.
Invariant Learning Robustness to out-of-distribution changes can be obtained by enforcing that the learning model is invariant to different environments or domains [2, 16, 44]. But there are many cases where we don’t have access to such labeled environments and we must discover them. Approaches like [9] partition the data into subsets that maximally contradict an invariant constraint, and apply algorithms for distributional robustness, like groupDRO [33], on those subsets, called environments.
Model guidance with explanations Most similar to our bias-based regularization is the use of model explanations to guide learning [31, 35, 27, 42]. Using differentiable attribution methods like Grad-CAM [34], they constrain the attention of the trained models to only fall on the class instances, segmented at the input level. A limitation of these approaches is that they can’t prevent the model from learning biases that are not spatially separated from the class in the input (e.g. color, shape, or accessories of the subject).
3 Our Method
For a standard classification task, we aim to identify biases acquired by a model through training on a new dataset. By dataset biases we refer to class-independent concepts, whose presence in the samples greatly affects the class label distribution. By concepts we refer to words or expressions with a semantic load, that are most relevant in describing a given context. For instance, to extract concepts from an image dataset, we begin with the image captions, extract keywords from them, and use the keywords as concepts. Through biases of a classifier $f_{\theta}$ we refer to concepts causally unrelated with a class, whose presence in the input significantly changes the distribution of class probabilities predicted by $f_{\theta}$. We further classify them as positively correlated biases w.r.t. to a class $k$ if the presence of these biases in the input increases the probability of $f_{\theta}$ predicting the class $k$ and negatively correlated biases if they decrease it.
Setup We start with a dataset with samples $(x_{j},y_{j})$, in which we search for biases, through the representations offered by a foundation model $M$, capable of embedding both the input samples $x_{j}$ and the concepts $c_{i}$ into the $\mathbb{R}^{D}$ vector space. We further assume that their cosine similarity $cos(M(c_{i}),M(x_{j}))$ is higher when the concept $c_{i}$ is present in sample $x_{j}$ than when it is not. We first construct a set of concepts $(c_{i})_{1\leq i\leq q}$ in textual form, that are present in the training data. To spot the biases among them, we train a linear layer on top of the input data representations given by $M$. Next, through concept filtering, we identify concepts that are causally unrelated to the classes, revealing our bias candidates. Finally, we rank them based on how strongly they influence the predictions of this linear classifier and automatically decide the bias score cut-off. In the following equations, we use the dot product instead of the cosine similarity, as all the vectors we use are normalized. The main steps of our method (also revealed in Fig. 1, and in more detail in the Appx. C algorithm), are the following:
Initialization We initialize $w_{k}$, the weights of class $k$ in the linear layer, with the embedding of its corresponding class name, extracted by the model $M$:
$w^{0}_{k}=M(\text{class\_name}_{k});\;\forall 1\leq k\leq|\mathcal{Y}|=N.$ (1)
Figure 2: The maximum distance between the reference line and the smoothed bias scores gives the threshold for our dynamic cut-off heuristic.
Step 1: Model training We train the weights using ERM [37] on our dataset of interest. Through learning, the class concepts (represented by the weights in the linear layer) naturally shift from the original initialization towards concepts that can effectively distinguish between class samples, given the current dataset, as visually presented in Fig. 1. This includes the classes themselves and their (positive) biases as well. At the end of the training, the weights can be expressed as a weighted sum of these concepts.
Step 2: Bias identification Using the learned weights, we continue to identify which concepts $c_{i}$ are either biases positively or negatively correlated with each class $k$. First, we filter out the class-related concepts. We then rank the remaining class-neutral concepts and select the top ones as biases based on a dynamic thresholding strategy, as follows:
a. Filter concepts We first aim to filter out concepts that are indeed related to a certain class, leaving only concepts that are causally unrelated to a class (we call them class-neutral concepts). These concepts are proper candidates for biases because they are not necessary for the recognition of a class (e.g. a forest background is well correlated with species of landbirds, but we want to prevent the model from using this causally unrelated correlation). The exact pipeline and methods are detailed in Sec. 4.
b. Rank class-neutral concepts For each class-neural concept $c_{i}$ and class $k$, we rank the concepts based on their similarity, using the following positive-bias score:
$s_{k,i}^{+}=w^{\top}_{k}M(c_{i})-{\underset{k^{\prime}=\overline{1,N}}{\text{min}}}w^{\top}_{k^{\prime}}M(c_{i}).$ (2)
Intuitively, we want concepts specific to one class but not to all the others. For this, for each class $k$, we select concepts highly correlated to it, while not correlated to at least one other class. For the negatively correlated biases, the score $s_{k,i}^{-}$ is similar, with the difference that we are taking into account the dissimilarity scores ($-w^{\top}_{k}M(c_{i})$).
Table 2: Validation procedure for three positively correlated biases found using ConceptDrift for CLIP ViT-L/14 fine-tuned on ImageNet. We evaluate the model’s capability to recognize a depicted (correct) class before and after the introduction of a biasing factor. For each prompt, $1000$ images are generated using FLUX.1-dev. We observe throughout all considered scenarios, a significant drop in the model’s capacity to identify the correct class when the biasing factor is involved. We also observe a meaningful increase in the likelihood of having the induced class predicted, even though it is not illustrated.
Exploited Bias $\rightarrow$ Avg. Softmax Activations (%) Samples Predicted As (%)
Induced Class Correct Class Prompt Correct Class Induced Class Correct Class Induced Class
firemen (bias) $\rightarrow$ fire truck (class) peafowl a photo of a peafowl 99.9 0.0 100.0 0.0
firemen and a peafowl 10.6 (-89.3 ) 78.9 (+78.9) 5.3 (-94.7) 93.4 (+93.4)
reading a newspaper (bias) $\rightarrow$ crossword (class) Mexican hairless dog a photo of a mexican hairless dog 45.2 0.0 47.5 0.0
a man reading a newspaper in a chair with a mexican hairless dog in its lap 2.0 (-43.2) 21.6 (+21.6) 0.9 (-46.6) 36.6 (+36.6)
shrimp (bias) $\rightarrow$ American lobster (class) Bernese Mountain Dog a photo of a Bernese Mountain Dog 90.2 0.0 99.8 0.0
shrimp and pasta near a Bernese Mountain Dog 8.1 (-82.1) 14.8 (+14.6) 10.6 (-89.2) 37.2 (+37.2)
c. Dynamic thresholding towards biases We select the biases of each class using a dynamic threshold for the bias scores above. We keep the highest-ranked ones, both for positive and negative bias scores. We remove the null scores and then apply a mean filter with a window size of $r$ to smooth the curve of scores. We denote the scores obtained at this step with $(\overline{s}_{k,i})_{1\leq i\leq p}$, with $p=q-r+1$. We then select the top $m$ ones, as positive biases for class $k$, where the $m$ index is defined as:
$m=\lfloor r/2\rfloor+\arg\max_{i}\left(\overline{s}_{k,1}-i\frac{\overline{s}_{k,1}-\overline{s}_{k,p}}{p-1}-\overline{s}_{k,i}\right).$ (3)
The intuition is that $m$ represents the index where the smoothed score curve deviates most from the straight reversed diagonal line, connecting $\overline{s}_{k,1}$ and $\overline{s}_{k,p}$ (see Fig 2).
Figure 3: Qualitative results depicting a series of biases uncovered employing our proposed ConceptDrift approach for CLIP ViT-L/14 fine-tuned on ImageNet [10]. All images are generated using FLUX.1-dev. In each scenario, a (correct) ImageNet class is depicted alongside a bias found to be positively correlated with a different (induced) class. Alongside each image, we provide the names of the correct and predicted classes, together with the softmax activations of their respective neurons. Though in most scenarios a single ImageNet class is clearly depicted, the model chooses to ignore it and predict a class that is not illustrated, clinging to the spurious correlations learned.
4 Experimental Setup
Foundation models (FM) We used mGTE (gte-large-en-v1.5 [45]) for text embeddings in CivilComments [5], and OpenAI CLIP ViT-L/14 [30] for text and images otherwise.
Concept extraction For the image classification task, we first use the GIT-Large  [41] captioning model (trained on MSCOCO [19]) to obtain descriptions of the dataset’s images. Next, we extract concepts from the captions (or directly from the text samples for the text classification task), using YAKE [7] keyword extractor, taking the top $256$ n-grams for $n=3,5$.
Concepts filtering We use Llama-3.1-8B-Instruct [11] to remove class instances from the concepts extracted at the previous step. We also apply post-processing using WordNet [24] to catch obvious class instances that the LLM might miss. To do this, for each class we specify a name used to search for synsets in the WordNet [24] hierarchy (e.g. bird for Waterbirds) and then remove individual words that match with any hyponym or hypernym of those synsets.
Training We train the linear layer on $L_{2}$ normalized embeddings extracted by these models using the PyTorch [26] AdamW [23] optimizer with a learning rate of $41e-$ 4, a weight decay of $51e-$ 5 and a batch size of $1024$. We use the cross entropy loss with balanced class weights as the objective. The weights of the layer are normalized after each update and we use CLIP’s [30] temperature to scale the logits. We use the class-balanced accuracy on the validation set as the criterion for model selection and early stopping.
4.1 Datasets
Waterbirds [40] is a common datasets for generalization and bias mitigation. It is created from CUB [39], by grouping species of birds into two categories, landbirds and waterbirds, each being associated with a spurious correlation regarding its background, land and water respectively.
CelebA [21] is a large-scale collection of celebrity images (over 200000), widely used in computer vision research. The setup for using it in a generalization context [22] consists of using the Blond_Hair attribute as the class label and the gender as the spurious feature.
CivilComments [5] is a large collection comprised of $1.8$ million online user comments. This dataset is used employed in NLP bias and fairness research concerning different social and ethnical groups.
ImageNet [10] is a larger-scale popular dataset for image classification ($1000$ classes, with approximately $1300$ training samples and $50$ validation samples per class).
5 Bias identification evaluation
For a proper quantitative evaluation of our proposed biases, beyond the subjectivity of the qualitative aspects, we combine the bias identified by ConceptDrift with different components: in Sec. 5.1 we generate samples exploiting the discovered biases; in Sec. 5.2 we assess GroupDRO [33] robustness over training set partitions based on the proposed biases groups; in Sec. 5.3 we evaluate bias-based training regularization; and in Sec. 5.4 we use a minimally bias-enhanced prompting for zero-shot prediction with a Foundational Model. An exhaustive list of the extracted biases for all datasets can be found in Appx. B.
5.1 ImageNet biases
We employ ConceptDrift to uncover a series of biases and spurious correlations plaguing the decision-making process of OpenAI’s CLIP ViT-L/14 fine-tuned on ImageNet. As the biases our method reveals lie beyond the restricted scope of the validation set, our sole recourse for illustrating and validating these findings at scale lies in generating synthetic images. To this end, we have employed a quantized version of FLUX.1-dev [17]. To validate the impact of the biases, we prompt the generative model to depict a chosen (correct) class alongside a non-ImageNet object that we identified as a positively correlated bias for a different class.
The validation process is presented for three distinct scenarios in Tab. 2. Each scenario is defined by a pair of classes: a correct class which is illustrated in the image and another class which is absent, but induced through the presence of a biasing object or activity. We measure the impact of the bias by means of comparing the model’s ability to predict the correct class before and after the introduction of the biasing factor. We generate $1000$ images for each prompt, and we evaluate the model both in terms of the predicted class and in terms of the activations of the class neurons. We observe throughout all considered scenarios, a significant drop in the model’s capacity to identify the correct class when the biasing factor is involved, and a significantly increased likelihood to have the induced class predicted, even though it is not illustrated in any way, shape or form in the generated image.
Within the same context, we present a series of qualitative examples in Fig. 3. We emphasize that, even though throughout most of these samples, a single ImageNet class is clearly depicted, the model chooses to ignore it and label the image as a completely different class, not illustrated at all in the image, solely based on the presence of a non-ImageNet object. As an example, the network chooses to completely ignore the peafowl, present right at the center of the second image, and labels the image as a ”Fire Truck” based on the presence of the firefighters.
The model at hand is generally considered to be a robust state-of-the-art model, benefiting from ample pre-training, achieving an $82.4\%$ top-1 accuracy on the ImageNet validation set. We emphasize through our experiment that even under these reassuring circumstances, critical reasoning flaws can make their way through in a production-ready model, undetected by validating and examining the model’s performance on held-out data (the traditional approach for identifying biases). It is through shifting the approach towards investigating the model’s weights that we can anticipate and prevent serious biases from making their way into the final version of the model.
Throughout this validation process, we have manually ensured that the generated data does not contain troubling artifacts. For instance, it is the case that even though the prompt only asks for a firefighter, in certain circumstances the generative network will also depict a fire truck alongside the other features requested by the prompt, against our desires. Additionally, sometimes the generative model is not precise and may depict a dog or a bird species different than the one we asked for. These examples of faulty generation are not exhaustive. It is therefore the case that prompts have to be carefully designed and images carefully inspected, in order to ensure the quality of the evaluation. Under these circumstances, we have endeavored to illustrate and validate but a small set of the remarkable flaws, ascribing a complete examination of our findings to future work. However, the complete list of ranked concepts for each class, determined automatically by employing our ConceptDrift approach, is available for consultation in Appx. B.
Table 3: Learning in the context of 100% spurious correlations. In the absence of samples with negatively correlated biases, GroupDRO performs similarly with the standard ERM. In contrast, bias regularization using our biases yields excellent outcomes in this setup, improving the worst group accuracy above the considered baselines: ERM, GroupDRO and the bias regularization with random causally-unrelated concepts (obtained after filtering in Step2a.).
Waterbirds (Acc % $\uparrow$) CelebA (Acc % $\uparrow$) CivilComments (Acc % $\uparrow$)
Method Worst Avg. Worst Avg. Worst Avg.
ERM 43.2$\pm$5.7 72.7$\pm$2.2 9.6$\pm$1.0 58.2$\pm$0.4 18.6$\pm$0.3 49.9$\pm$0.2
GroupDRO 38.9$\pm$5.4 71.2$\pm$2.0 8.1$\pm$0.3 60.3$\pm$1.0 18.7$\pm$0.4 50.2$\pm$0.5
Bias regularization w/ random concepts (class-neutral) 46.6$\pm$2.7 75.3$\pm$1.1 9.4$\pm$0.0 61.4$\pm$2.0 19.1$\pm$1.6 50.8$\pm$0.9
Bias regularization w/ ConceptDrift biases (ours) 57.9$\pm$0.3 79.8$\pm$0.1 10.4$\pm$0.5 62.0$\pm$1.8 31.3$\pm$0.7 57.5$\pm$0.4
5.2 From biases to groups, towards generalization
Next, we move to the subpopulation shift setup [33] (where different biases determine subgroups in the data) and follow the evaluation protocol in SubpopBench [43]. Normally, ERM training on such data behaves poorly on minority groups, but methods like GroupDRO [33] can recover the performance, when the subgroups are given. We show that the biases discovered by our method can be used to create such subgroups, proving that they are indeed spuriously correlated concepts that a robust model should not rely on.
Table 4: Generalization task, evaluated by comparing the performance of an ImageNet-pretrained ResNet-50 model, trained with GroupDRO on dataset partitions that are based on the identified biases. ConceptDrift biases are competitive with SoTA.
. Waterbirds (Acc % $\uparrow$) CelebA (Acc % $\uparrow$) Training Bias identification Worst Avg. Worst Avg. ERM - 70.0$\pm$2.4 84.9$\pm$0.9 54.2$\pm$3.6 95.4$\pm$0.3 GroupDRO w/ B2T [15] partitioning into groups B2T [15] 83.2±$\pm 3.6$ 90.6±$\pm 2.0$ 87.6±$\pm 2.0$ 93.7±$\pm 0.4$ Lg [47] 81.4±$\pm 3.4$ 89.5±$\pm 1.3$ 89.3±$\pm 1.8$ 93.2±$\pm 0.6$ ConceptDrift (ours) 81.9±$\pm 2.2$ 90.0±$\pm 2.3$ 88.4±$\pm 1.4$ 93.6±$\pm 0.4$ Ground-truth groups 85.8±$\pm 2.5$ 90.1±$\pm 1.6$ 88.8±$\pm 2.0$ 93.4±$\pm 0.7$
For the experiments in Tab. 4, we use the following pipeline. Based on the identified biases, we split the training data into subgroups with the partitioning technique from B2T [15]. Next, we train the model with GroupDRO, constraining it to attend equitably to the subgroups. If the concepts represent biases that will lead the model to learn spurious correlations through ERM, then under GroupDRO training, the model will reduce its reliance on them, surpassing the performance of ERM on the minority subgroups.
Tab. 4 shows that our identified ConceptDrift-biases are indeed useful for GroupDRO, far surpassing the ERM baseline and obtaining results that are competitive with the other bias identification methods in terms of worst-group performance on the Waterbirds and CelebA datasets. Note that the partitioning method was not tuned to accommodate our biases (e.g. our biases include actions and attributes, while the other methods only propose nouns).
Loss correlates with the presence of biases In the setup described above, we look at the correlations between the loss values and the bias-to-sample similarities. We compare ERM with GroupDRO applied on subgroups obtained with our biases and the B2T partitioning strategy. See in Appx. D how after GroupDRO with our proposed biases, the correlations significantly decrease, revealing that the model is less prone to make mistakes on the biased samples. Intuitively, this also indicates that the model is less reliant on the biases.
5.3 Biases used for training regularization
We previously showed in Sec. 5.2 that partitioning datasets using qualitative biases, followed by GroupDRO training, can significantly improve the robustness of a model to the bias-induced distribution shifts. We now investigate a scenario where prior methods for robustness like GropuDRO or others [33, 44, 14] are not applicable. In a setup with 100% bias-to-class correlations, the absence of the diverse bias-class associations renders these existing methods ineffective. We thus propose the task of learning in this perfect correlated setup, removing all bias conflicting samples from the training and validation sets.
We use the discovered biases to impose regularization upon the trained models, in particular, regularization for linear probing, learned on top of frozen CLIP embeddings. Intuitively, we constraint each bias to be equally distanced from the class weights. We formulate this as a MSE between the similarity of $b$ with a class weight and the average similarity of $b$ to all the class weights, scaled by $\tau$ (CLIP’s temperature):
$\mathcal{L}_{reg}(b)=\frac{\tau^{2}}{N}\sum^{N}_{k=1}\left[w^{\top}_{k}M(b)-sg\left(\frac{1}{N}\sum^{N}_{j=1}w^{\top}_{j}M(b)\right)\right]^{2},$ (4)
with sg being the stop gradient operator. The final loss is
$\mathcal{L}=\mathcal{L}_{ERM}+\alpha\frac{1}{\left|\mathcal{B}\right|}\sum_{b\in\mathcal{B}}\mathcal{L}_{reg}(b),$ (5)
where $\mathcal{B}$ is the set of selected biases and $\alpha=0.1$.
In Tab. 3, we present the results of linear probing with $\mathcal{L}$ loss, in this perfect spurious correlation scenario. The improvement in worst group accuracy shows that the learned classification weights are less reliant on the biases, thus more robust and better aligned with the classification task.
Table 5: Results for the Foundation Model (FM) Zero-shot prompting task. We help the model account for biases by explicitly introducing them in the prompting scheme. We perform tests employing biases discovered by different methods and evaluate the zero-shot performance of the FM. Aided by the biases discovered through our proposed method, the model’s performance, measured in terms of worst group accuracy, is significantly enhanced across all test scenarios. Our method surpasses in this regard all previous bias-identification approaches.
Waterbirds (Acc % $\uparrow$) CelebA (Acc % $\uparrow$) CivilComm (Acc % $\uparrow$)
Method Worst Avg. Worst Avg. Worst Avg.
FM zero-shot 35.2 90.7 72.8 87.4 33.1 83.4
FM w B2T [15] 48.1 86.1 72.8 88.0 - -
FM w SpLiCE [4] - - 67.2 90.2 - -
FM w Lg [47] 46.1 85.9 50.6 87.2 - -
FM w ConceptDrift 50.3 86.3 73.1 85.7 53.2 71.0
Table 6: Ablation. Following the zero-shot bias-augmented prompting setup, we variate the cut-off threshold for considering causally-unrelated concepts as biases and also try to address the text-image modality gap.
Waterbirds (Acc % $\uparrow$) CelebA (Acc % $\uparrow$) CivilComm (Acc % $\uparrow$)
Variations Worst Avg. Worst Avg. Worst Avg.
top-30 candidates 46.3 86.1 66.2 86.4 46.9 71.1
top-20% candidates 46.1 86.0 64.8 86.7 48.8 66.2
modality gap: closed 48.7 85.9 72.7 85.3 - -
ConceptDrift
dynamic threshold
modality gap: open 50.3 86.3 73.1 85.7 53.2 71.0
5.4 Bias-aware zero-shot prompting
To further validate our identified biases, we follow Kim et al. [15] and evaluate them in the context of zero-shot classification task. We augment the initial, class-only, related prompt with the bias only through a minimal intervention (e.g.’a photo of a {cls} in the {bias}’ (see Appx. A). For each class we create a prompt with each identified bias. When classifying an image, we take into account only the highest similarity among the prompts of a class (zero-shot with max-pooling over prompts). The results in Tab. 5 show how the biases automatically selected by our method improve the worst group accuracy over the initial zero-shot baseline and other state-of-the-art bias-extracting solutions in all the tested datasets. This highlights the relevance of the biases automatically extracted by ConceptDrift.
Ablations We validate several key decisions in our algorithm in Tab. 6, using the zero-shot classification task and the biased-enhanced prompting. We notice that the number of chosen class-unrelated concepts is important, as taking too many of them adds noise to the prompts and lowers the performance. Nevertheless, dynamically choosing the threshold, as described in Sec. 3-Step2c., proves to be a good strategy for adapting the cut-off across datasets. Following prior observations regarding the modalities gap between text and image embedding space [18, 46], we follow Liang et al. [18] and subtract half of the gap from the embeddings, then re-normalize them, ending in marginally lower performance w.r.t. not addressing the modality gap.
6 Conclusions
We introduce ConceptDrift, a method that identifies biases through a weight-space and filtration approach, that can be scaled to large datasets. By leveraging CLIP models with aligned image-text embedding, we can easily identify concepts that significantly influence class predictions, while filtering undesirable concepts. We apply this method to multiple modalities on datasets like Waterbirds, CelebA, CivilComments, and ImageNet. We validate that the desired biases are useful by employing them in different methods for improved generalization. This includes a) creating subgroups that promote robust generalization b) a novel bias regularization that works even under $100\%$ correlation of biases and classes, where previous methods fail c) using better bias-aware zero-shot promoting. Using ConceptDrift we identify biases in ImageNet and generate synthetic images that are able to fool strong pre-trained methods, validating the relevance of the identified biases in larger datasets. This method opens the door for automatically finding biases in large settings without explicit human intervention.
Supplementary Material
Appendix A Zero-Shot Prompts
In Tab. A.1, we structured the class names used for initializing the initial weights of the linear layer, along with the prompt templates employed in the zero-shot classification experiments discussed in Sec. 5.4.
Table A.1: Class names and prompts used in the zero-shot classification task.
Waterbirds CelebA CivilComments
class names waterbird landbird non-blonde hair blonde hair non-offensive offensive
zero-shot prompt a photo of a {cls} a photo of a person with {cls} {cls}
bias-enhanced prompt a photo of a {cls} in the {bias} a photo of a {bias} with {cls} a/an {cls} comment about {bias}
Appendix B Biases and top class-neutral concepts
We present the exhaustive list of biases for the Waterbirds (Tab. B.1 & B.2), CelebA (Tab. B.3 & B.4) and CivilComments (Tab. B.5 & B.6) datasets. We also present top class-neutral concepts for ImageNet classes. Concept filtering on ImageNet was performed using WordNet relationships alone, without the intervention of a Large Language Model. Accounting for the size of the dataset, we will publish the available data on our repository upon acceptance, and we will restrict the presentation within the context of the current format to a few classes for illustrative purposes in Tab. B.7 -B.15.
Table B.1: Top Waterbirds class-neutral concepts for ”landbird”.
Landbird Score
forest floor 0.055562317
forest next to a tree 0.053587496
bamboo forest floor 0.05134508
forest of bamboo 0.04781133
forest 0.047080815
snowy forest 0.044688106
ground 0.043406844
field 0.043168187
log 0.043052554
standing on a forest floor 0.041143
grass covered 0.040526748
tree branch in a forest 0.039670765
forest with trees 0.03949821
tree in a forest 0.039123535
bamboo forest 0.03876221
front of bamboo 0.037381053
mountain 0.036155403
forest of trees 0.03600967
flying through a forest 0.035692394
platform 0.03565806
standing in a forest 0.034528017
hill 0.03341371
Table B.2: Top Waterbirds class-neutral concepts for ”waterbird”.
Waterbird Score
swimming in the water 0.11482495
water lily 0.10905403
boat in the water 0.1066975
floating in the water 0.106155455
water 0.106134474
flying over the water 0.10561061
standing in the water 0.10444009
sitting in the water 0.103776515
body of water 0.09977633
water in front 0.0902465
standing in water 0.086544394
water and one 0.08424729
swimming 0.07818574
standing on a lake 0.06565446
flying over the ocean 0.06509364
flying over a pond 0.06463468
boats 0.061231434
lifeguard 0.06122452
flying over a lake 0.060126305
boat 0.0571931
pond 0.053261578
Table B.3: Top CelebA class-neutral concepts for ”non-blonde”.
Non-Blonde Score
hat on and a blue 0.13952243
hat on and a man 0.13853341
man in the hat 0.13803285
man who made 0.13307464
man behind 0.13247031
man with the hat 0.13186401
man is getting 0.12861347
actor 0.12726557
dark 0.12713176
man in a blue 0.1269682
person 0.12541258
man in the blue 0.124844134
man is not a man 0.12308431
man 0.12290484
large 0.1228559
shirt on in a dark 0.12170941
hat 0.121646166
close 0.12146461
man with the blue 0.12136656
man face 0.12130207
Table B.4: Top CelebA class-neutral concepts for ”blonde”.
Blonde Score
yellow -0.021363914
white -0.04120779
golden -0.041870475
woman with blowing -0.042802572
woman with laughing -0.046326935
girl with blue eyes -0.050636113
woman with playing -0.0510059
girl next door -0.05101496
woman with green -0.05205536
sweater -0.054968417
gold -0.057154775
scarf -0.0585289
pale -0.05990547
woman with blue eyes -0.05999875
tank top -0.060728252
Table B.5: Top CivilComments class-neutral concepts for ”non-offensive”.
Non-offensive Score
allowing 0.07341421
work 0.06982881
made 0.069063246
talk 0.06858361
none are needed 0.067236125
check 0.06664443
helping keep the present 0.06531584
policy 0.06339955
campaign 0.06333798
involved in the first place 0.063222766
Cottage 0.063149124
IDEA 0.06310266
stories 0.0625782
job 0.06236595
allowed 0.062137783
latest news about the origin 0.062061936
giving others who have experienced 0.061925888
proposed 0.061897278
one purpose 0.06122935
starting 0.061154723
small 0.061071455
question 0.060854554
practice 0.060740173
raised 0.060681045
entering 0.060585797
registered 0.060475767
beliefs 0.060165346
accept that they are promoting 0.060070753
Security 0.059328556
new 0.059324086
subject 0.058983028
close 0.058632135
views 0.058573127
Hold 0.058341324
reality for a change 0.058261245
built at that parish 0.057885766
rest 0.057804525
historic 0.057656527
concept 0.057422698
people 0.057151675
passage seems to in reflection 0.05699992
attempt 0.056797385
Table B.6: Top CivilComments class-neutral concepts for ”offensive”.
”Offensive” – Top Concepts Score
hypocrisy 0.046756804
troll 0.035944045
silly 0.029536605
hate 0.013704538
silly how do you study 0.00325954
spite 0.002645433
kid you have the absolute 0.001619577
die on the golf -0.001492977
Muslim Canadians Party -0.002062559
serve them at the lunch -0.00358063
culture is no worse -0.005233765
banned -0.006861687
pretend they are all goodness -0.007072449
pop that particular rainbow -0.008466244
member becomes so sick -0.008648455
men and she wanted -0.008733362
Canadian Muslims -0.009060442
destroy -0.010027111
six Baltic -0.01069364
customers when it not acceptable -0.011366904
Table B.7: Top ImageNet class-neutral concepts for ”American Lobster”.
”American Lobster” – Top Concepts Score
pasta with shrimp 0.3148532
adirondack sitting 0.31178916
large shrimp 0.3085378
shrimp and pasta 0.30821544
shrimp cooking 0.3032636
large cast cooking 0.30282205
pasta with shrimp and cheese 0.2984723
adirondack 0.29169592
and mussels 0.2904386
legs and other seafood 0.28870153
close up of a shrimp 0.2871693
of pasta with shrimp 0.2823377
seafood 0.2808096
mussels 0.28018713
cast cooking 0.27951774
shrimp and cheese 0.27841187
shrimp 0.2726224
Table B.8: Top ImageNet class-neutral concepts for ”Bald Eagle”.
”Bald Eagle” – Top Concepts Score
osprey flying 0.2888234
osprey 0.27377927
row of american flags 0.26513425
group of american flags flying 0.25681192
american flag and american flag 0.23945728
emu standing 0.23752406
yellowstone national 0.23740456
Table B.9: Top ImageNet class-neutral concepts for ”Ballpoint Pen”.
”Ballpoint Pen” – Top Concepts Score
wearing a pilot 0.36252645
markers 0.36050144
sharpie 0.34563553
stylus 0.34399948
pair of eyeglasses 0.32426757
notepad 0.3228797
calligraphy 0.3222967
paint and markers 0.32198787
close up of a needle 0.32131955
on a straw 0.3209229
and markers 0.32090995
eyeglasses 0.31017447
dots 0.30764964
crayons 0.30598855
Table B.10: Top ImageNet class-neutral concepts for ”Crossword”.
”Crossword” – Top Concepts Score
reading a newspaper 0.30469692
man reading a newspaper 0.29045385
crochet squares 0.28316277
sitting on a newspaper 0.27749887
newspaper sitting 0.27106437
crochet squares in a square 0.2708223
newspaper that has the words 0.26934764
holding a newspaper 0.26375395
newspaper while sitting 0.26288068
newspaper laying 0.25538272
square with a few crochet 0.2461972
square with a crochet 0.24225119
square of crochet squares 0.23986068
checkerboard 0.23797607
crochet blanket in a square 0.2364017
square of square crochet 0.23590976
on a newspaper 0.2357213
crochet square with a crochet 0.23569846
crochet blanket with a crochet 0.23438567
newspaper 0.23315597
crochet with a square 0.23304509
crochet square 0.23259673
newspaper sitting on 0.23098715
square of crochet yarn 0.2291883
square crochet 0.22903368
crochet blanket with a square 0.22888878
crochet square sitting 0.22860557
crochet in a square 0.22856355
square with a single crochet 0.22752959
free crochet 0.22748157
newspaper with 0.22672665
is on a newspaper 0.2257084
crochet blanket 0.22547376
checkered blanket 0.22514643
square of crochet 0.22324148
Table B.11: Top ImageNet class-neutral concepts for ”Coffeemaker”.
”Coffeemaker” – Top Concepts Score
kettle sitting on 0.43714887
with a kettle 0.4207235
thermos 0.40833473
kettle 0.40526068
kettle sitting 0.3881535
stovetop maker 0.37747166
kettle kettle 0.37671012
kettle kettle kettle kettle 0.37649006
vases and vases 0.37567452
kettle kettle kettle 0.37547356
flask 0.37485123
kettle kettle kettle kettle kettle 0.37305972
large canister 0.37186915
flasks 0.36940324
decorative vases sitting 0.3683877
kitchen aid 0.36793774
milkshakes 0.36605325
large pottery 0.36506
set of kitchen 0.3650242
cookbook 0.36469316
vases sitting 0.3643943
Table B.12: Top ImageNet class-neutral concepts for ”Doormat”.
”Doormat” – Top Concepts Score
brick sidewalk 0.37059835
laying on gravel 0.34978455
laying on a carpeted 0.34279323
crochet blanket 0.3421431
sitting on a carpeted 0.34104648
laying on a step 0.3400078
crochet blanket in a square 0.33642814
brick walkway 0.3319164
carpeted floor 0.33067068
on a brick sidewalk 0.32878387
crochet blanket with a crochet 0.32599914
floor with a welcome 0.32548892
square of crochet yarn 0.32428077
crochet blanket made 0.32336423
carpeted staircase 0.3227385
dot blanket 0.3213501
mosaic floor 0.32125634
crocheted blanket 0.31886423
on a blanket 0.3182252
crochet blanket with a square 0.31753486
square of crochet squares 0.3169018
crochet squares 0.3152317
standing in a doorway 0.3126963
Table B.13: Top ImageNet class-neutral concepts for ”Eraser”.
”Eraser” – Top Concepts Score
crayons 0.38733196
chalk 0.37349075
graphite 0.36778685
crayon 0.36178917
charcoal 0.351962
lip balm lip 0.35033816
band aid cookie 0.34825876
lip balm 0.3409983
nose sticking 0.34016216
sticking 0.33971623
markers 0.33940658
matchbox 0.33480892
wand 0.33415005
stylus 0.3318595
band aid card 0.33174193
toothbrush 0.33009088
office supplies 0.32956824
band aid flexible 0.32946587
Table B.14: Top ImageNet class-neutral concepts for ”Fire Truck”.
”Fire Truck” – Top Concepts Score
firefighter spraying 0.34155905
group of firefighters 0.3347583
firefighters 0.33051446
firefighter 0.31450543
firefighter wearing 0.2843979
hydrant spraying 0.26793447
firefighter wearing a 0.26590723
firefighter cuts 0.2613345
farmall parked 0.2591076
dashboard with flames 0.25825307
flames painted 0.2540929
Table B.15: Top ImageNet class-neutral concepts for ”Guacamole”.
”Guacamole” – Top Concepts Score
tomatoes and avocado 0.33948907
avocado 0.3125001
nachos 0.30761188
ham and parsley 0.2960047
with avocado 0.2952027
colorful mexican 0.28313732
bacon and parsley 0.27773544
of avocado 0.2754345
side of salsa 0.27470407
peridot 0.2734925
salsa 0.27124816
nachos with 0.27092364
lime body 0.27088284
cheese and parsley 0.26951405
pile of limes 0.26944226
peas and bacon 0.2671204
limes and limes 0.26682717
lime cut 0.2662533
nachos with and 0.26580203
pasta with peas 0.2626204
tacos 0.2619322
pasta with ham and parsley 0.26183394
tomatoes and cilantro 0.26161948
Appendix C Algorithm
We present the formal implementation of our proposed ConceptDrift approach in Alg.  1. We annotate the main steps presented in Sec. 3. At line 2, we initialize the class weights of our linear probing layer with the text embeddings of the class names, i.e. the zero-shot classification weights of the foundation model $M$. We then fine-tune the layer, at line 3, on the given dataset. At lines 5 and 6, we filter the list of concepts and compute the embeddings for the remaining class-neutral concepts. The filtering can be performed by means of employing WordNet associations, Large Language Models, or both. With the embeddings of the class-neutral concepts at hand, we proceed to determine the biases for each class. At line 7, we initialize our set of biases with an empty set, and we compute the similarities between each selected concept and each class at line 8.
We further proceed to iterate through each class (line 9). For a given class $k$, we compute the set of scores for all concepts with respect to class $k$ and store it in $s_{k}$, at line 11. For each class-neutral concept, its ranking score for the current class is the difference between its similarity to the current class and the smallest similarity with a different class. Lines 12-14 formally implement the dynamic thresholding procedure proposed in Sec. 3. Finally, we select the concepts with ranking scores above the computed threshold and store them as biases for class $k$, at lines 15 and 16.
Algorithm 1 ConceptDrift
Input:
M - Foundation Model with associated text encoder
$(\mathcal{X},\mathcal{Y})$ Training set
$(\mathcal{X}_{val},\mathcal{Y}_{val})$ Validation set
K - list of class names
$C_{all}$ list of all concepts
r - window size for dynamic thresholding smoothing
Output: Identified positively correlated biases $\mathcal{B}$.
1:    // Step 1. Model Training
2:   $\mathbf{W^{0}}\leftarrow M(K)$
3:   $\mathbf{W}\leftarrow\mbox{ERM}\left(\mathbf{W^{0}},\mbox{M},(\mathcal{X},\mathcal{Y}),(\mathcal{X}_{val},\mathcal{Y}_{val})\right)$
4:    // Step 2. Bias identification
5:   $C\leftarrow$Filter$(C_{all})$ // 2a. Filter concepts: LLM/WordNet
6:   $\mathbf{C}^{*}\leftarrow\mbox{M}(C)$
7:   $\mathcal{B}\leftarrow\emptyset$
8:   $\mathbf{S}\leftarrow\mathbf{W}^{\top}\mathbf{C}^{*}$ // 2b. Rank class-neutral concepts
9:   for $k\in\overline{1,|K|}$ do
10:       // 2c. Dynamic thresholding, for each class $k$
[…]
12:      $\overline{s}_{k}=\mbox{mean\_pool}(\mbox{reversed}(\mbox{sorted}(s_{k})),r)$
13:      $p=|C|-r+1$
14:      $th_{k}=\lfloor r/2\rfloor+\arg\max_{i}\left(\overline{s}_{k,1}-i\frac{\overline{s}_{k,1}-\overline{s}_{k,p}}{p-1}-\overline{s}_{k,i}\right)$
15:      $b_{k}=\mathcal{C}\left[s_{k}\geq th_{k}\right]$
16:      $\mathcal{B}=\mathcal{B}\cup(k,b_{k})$ // Positively correlated biases
17:   end for
18:   return  $\mathcal{B}$
Appendix D Loss to bias similarity correlation
To validate that the environments discovered through the partitioning approach from B2T [15], using the biases uncovered with our method, are indeed effective in reducing the model’s reliance on shortcuts, we analyzed the correlation between the similarity of each training image to biases and its training loss value under both ERM and GroupDRO training, using our identified groups. The results show a reduction in correlation scores across all biases, demonstrating that the revealed groups are relevant to the dataset’s underlying distribution, and can be effectively utilized with specific algorithms to mitigate the model’s dependence on spurious correlations. Fig. D.1 shows the Pearson correlation scores after an epoch of training on Waterbirds, on a subset of all biases.
Figure D.1: Correlation(sample_loss, sample_to_bias similarity) under ERM/GDRO after one epoch of training on Waterbirds. Loss correlation w/ biases, ERM vs GroupDRO using groups created with the B2T partitioning method. It can be seen that, when training with ERM, loss value is highly correlated with the biases. In contrast, GroupDRO reduces the correlations, intuitively showing that biases discovered with our method are closely related to the ground truth groups of the dataset, being used as shortcuts by the model unless mitigated.

### v3

@@
-ConceptDrift: Uncovering Biases through the Lens of Foundation Models
+WASP: A Weight-Space Approach to Detecting Learned Spuriousness
 Abstract
-An important goal of ML research is to identify and mitigate unwanted biases intrinsic to datasets and already incorporated into pre-trained models.
-Previous approaches have identified biases using highly curated validation subsets, that require human knowledge to create in the first place.
-This limits the ability to automate the discovery of unknown biases in new datasets.
-We solve this by using interpretable vision-language models, combined with a filtration method using LLMs and known concept hierarchies.
-More exactly, for a dataset, we use pre-trained CLIP models that have an associated embedding for each class and see how it drifts through learning towards embeddings that disclose hidden biases.
-We call this approach ConceptDrift and show that it can be scaled to automatically identify biases in datasets like ImageNet without human prior knowledge.
-We propose two bias identification evaluation protocols to fill the gap in the previous work and show that our method significantly improves over SoTA methods, both using our protocol and classical evaluations.
-Alongside validating the identified biases, we also show that they can be leveraged to improve the performance of different methods.
-Our method is not bounded to a single modality, and we empirically validate it both on image (Waterbirds, CelebA, ImageNet), and text datasets (CivilComments).
+It is of crucial importance to train machine learning models such that they clearly understand what defines each class in a given task.
+Though there is a sum of works dedicated to identifying the spurious correlations featured by a dataset that may impact the model’s understanding of the classes, all current approaches rely solely on data or error analysis.
+That is, they cannot point out spurious correlations learned by the model that are not already pointed out by the counterexamples featured in the validation or training sets.
+We propose a method that transcends this limitation, switching the focus from analyzing a model’s predictions to analyzing the model’s weights, the mechanism behind the making of the decisions, which proves to be more insightful.
+Our proposed Weight-space Approach to detecting Spuriousness (WASP) relies on analyzing the weights of foundation models as they drift towards capturing various (spurious) correlations while being fine-tuned on a given dataset.
+We demonstrate that different from previous works, our method (i) can expose spurious correlations featured by a dataset even when they are not exposed by training or validation counterexamples, (ii) it works for multiple modalities such as image and text, and (iii) it can uncover previously untapped spurious correlations learned by ImageNet-1k classifiers.
+The code is available at https://github.com/bit-ml/WASP.
+spurious correlation, ImageNet, waterbirds, CelebA, CivilComments, CLIP, foundation models, subpopulation shift
+[…]
 1 Introduction
-Deep neural networks, and especially fine-tuned versions of foundation models, are commonly deployed in critical areas such as healthcare, finance, and criminal justice, where biased predictions can have significant societal consequences [1].
-Despite their impact, these models are often employed in their natural black-box state, i.e.
-as highly non-linear, multi-layered decision processes, lacking transparency or interpretability.
-Even if the pretrained model has been validated by the community, the dataset leveraged in the fine-tuning process can, and usually does, imprint the model with new biases.
-This issue is particularly concerning as biases can lead to undesired outcomes [6], reinforcing existing inequalities or creating new forms of discrimination.
-Subpopulation shift setups [40, 21, 5], featuring naturally occurring spurious correlations, provided researchers with a controlled environment for studying bias detection and prevention.
-Within this context, recent efforts began employing foundation models [15, 47] in their investigation protocols.
-However, current methods limit themselves to data analysis alone.
-For instance, the method proposed by Kim et al.
-[15] relies on validating the presence of a given object within the set of mistakes and its absence from within the set of the correctly classified samples within the validation set, in order to label it as a bias.
-Thus, data-restricted methods rely on the critical assumption that the validation set is exhaustive in some sense, that it contains samples capable of exposing all biases.
-This assumption is justified only within the context of a controlled experimental environment, such as having a subpopulation shift setup where all sources of bias are given or known.
-Any other real-world scenario hardly guarantees an exhaustive validation set.
-Figure 1: Following ConceptDrift’s steps for the ImageNet ”Fire Truck” class.
-In Step 1, during training, the classification weights $W$ drift from the initial class concept embedding $W^{0}$, outside the scope of relevant concepts, towards biases.
-In Step 2, our method filters out class-related concepts and, using an embedding-space scoring system, ranks and automatically selects the highest-ranking class-neutral concepts as biases.
-Leveraging these identified biases, we generate synthetic images that are able to fool even strong pre-trained methods.
-We expand the current usage of foundation models beyond the restricted scope of data analysis and propose a new direction for bias identification.
-Our method focuses on investigating the skewness of the model’s weights towards detecting and prioritizing spurious features as part of the decision-making process of the investigated model.
-We hereby propose a novel protocol for uncovering biases using foundation models such as CLIP [30] and mGTE [45], leveraging the topology of their embedding space to identify and name biases instilled by linear probing.
-Our protocol, dubbed ConceptDrift, is illustrated in Fig.
-1.
-We showcase how, during training, the weights of the final classification layer drift away from the textual representation of their associated class, towards representations of spurious attributes.
-We propose a ranking system based on embedding-space arithmetic to extract concepts that factor in the activation of class neurons and delineate the highest-ranking concepts that lie outside the semantic scope of the classes, as biases.
-Our main contributions are summarized as follows:
+Deep neural networks, and especially fine-tuned versions of foundation models, are commonly deployed in critical areas such as healthcare, finance, and criminal justice, where decisions based on spurious correlations (SCs) can have significant societal consequences (Angwin, Julia and Larson, Jeff and Mattu, Surya and Kirchner, Lauren, 2016; Caliskan et al., 2017).
+Even if the pretrained model has been validated by the community, the dataset leveraged in the fine-tuning process can, and usually does, imprint the model with new SCs.
+We present such an example in Fig.
+1, where a CLIP model fine-tuned on ImageNet-1k labels the image as a ”Fire Truck” even in the complete absence of any fire truck or truck-related features and even given the presence of a peafowl (which is an ImageNet-1k class) right in front and center.
+This is due to the fact that, as our proposed WASP protocol discovered, the model spuriously correlates firefighters (which are not an ImageNet-1K class) with the ”Fire Truck” class.
+Figure 1: Leveraging WASP’s identified spurious correlations, we generate synthetic images that fool strong pre-trained models.
+Subpopulation shift setups (Wah et al., 2011b; Liu et al., 2015a; Borkan et al., 2019), featuring naturally occurring SCs, provide researchers with a controlled environment for studying SC detection and prevention.
+Within this context, recent efforts have begun to employ foundation models (Kim et al., 2024; Zhao et al., 2024; Bhalla et al., 2024) in their investigation protocols.
+Some of these efforts (Bhalla et al., 2024; Zhao et al., 2024) focused on finding SCs through data analysis, without referencing model predictions.
+While these findings are valuable, they may not always be relevant.
+For example, an SC may not impact a machine learning model if the correlation is harder to learn than the actual class itself.
+To this end, efforts such as that of (Kim et al., 2024) targeted learned spurious correlations, focusing on investigating validation samples that the model misclassifies.
+However, this error-based approach assumes that the validation set is in some sense exhaustive and that the validation data are already able to expose the SCs learned by the model, which may be true for subpopulation shift setups, but hardly guaranteed in any other general setup.
+Figure 2: Following WASP’s steps for the ImageNet-1k ”Fire Truck” class.
+In Step 1, during training, the classification weights $W$ drift from the initial class concept embedding $W^{0}$, outside the scope of relevant concepts, towards spuriously correlated ones.
+In Step 2, our method filters out class-related concepts and, using an embedding-space scoring system, ranks and automatically marks the highest-ranking class-neutral concepts as SCs.
+Different from prior work, we aim to overcome the limitations of data and error analysis, revealing spurious correlations learned by machine models even in scenarios in which the training and validation sets do not present counterexamples able to expose the model’s flaws.
+To this end we propose WASP, a Weight-space Approach to detecting learned Spurious correlations.
+Our overall framework is illustrated in Fig.
+2 and further detailed in Sec.
+3.
+We leverage the topology of foundation models such as CLIP and mGTE and observe that during the process of training, the weights of the final classification layer drift away from the textual representation of their associated class, towards identifying and prioritizing representations of spurious attributes.
+We propose a ranking system based on embedding-space arithmetic to extract concepts that factor in the activation of class neurons and delineate the highest-ranking concepts that lie outside the semantic scope of the classes, as spurious correlation.
+Our work can be summarized as follows:
   1. 1.
-    We introduce ConceptDrift, a method capable of pin-pointing concepts that are relevant for the decision-making process of a model.
-We are the first to propose a weight-space approach for identifying biases that a foundation model learns from a given dataset, diverging from the current techniques oriented on validation set errors.
+    We introduce WASP, a weight-space approach to detecting learned spuriousness, offering a departure from the current error or data analysis methodology.
   2. 2.
-    We propose two novel evaluation strategies for the bias identification task.
-The first one evaluates the performance decay over generated samples that exploit the identified biases, highlighting the potential harm biases can cause to models.
-The second one assesses the performance gains achieved through bias-based regularization in a training setup with a 100% spurious correlation between classes and biases, focusing on mitigating biases and enhancing model robustness.
+    We show that our approach surpasses all state-of-the-art methods for identifying and naming spurious correlations in a given dataset.
+First, in its capacity to enhance the robustness of zero-shot models, and second, in its applicability to scenarios lacking spurious correlation counterexamples.
+Furthermore, in addition to evaluating our method on the established image datasets (Waterbirds, CelebA), we also validate its effectiveness on textual data (CivilComments).
   3. 3.
-    We validate ConceptDrift, using both established evaluation protocols and our newly introduced strategies.
-We achieve significant advances upon the SoTA bias identification techniques on classical image datasets: Waterbirds [40] and CelebA [21], and further extend the evaluation by including a text dataset: CivilComments [5].
-We also quantitatively confirm biases identified in ImageNet with respect to a foundation model.
-  4. 4.
-    We show that more robust methods can be obtained by devising constraints around the discovered biases. This includes subgroup constraints, training regularization, and zero-shot prompting.
-Table 1: Comparison for Bias Identification approaches, based on fundamental differences in methods.
-Method Root of bias identification Scoring bias candidates Fully auto Evaluation on large datasets Evaluation protocols employed
-SpLiCE [4] disentangled representations Lasso solver Masking concepts from decomposition
-B2T [15] mistakes in valid-set keywords frequency GroupDRO over bias-based groups
-Lg [47] class-specificity score latent-space arithmetic GroupDRO over bias-based groups Train on bias-exploiting samples
-ConceptDrift (ours) biases’ similarity to class weights latent-space arithmetic partially Validate on bias-exploiting samples (Sec.
-5.1) GroupDRO over bias-based groups (Sec.
-5.2) Bias-based training regularization (Sec.
-5.3)
-2 Related Work
-Bias identification for generalization Machine learning methods naturally capture relevant factors needed to solve a task.
-Nevertheless, many times, models also capture shortcuts [12], that are helpful in solving the task, but are not fundamental or essential for it.
-These shortcuts represent spurious correlations or biases, that don’t always hold, and should not be used for reliable generalization outside the training distribution, often leading to degraded performance [29, 3, 13].
-Prior works [15, 47, 4] have thus focused on identifying dataset biases, mainly through data analysis procedures.
-Kim et al.
-[15] identify concepts that are more prevalent in the misclassified examples in comparison with the correctly classified ones, while Zhao et al.
-[47] look for concepts that are specific to one class.
-In contrast, Bhalla et al.
-[4] identifies biases by training a linear layer on top of a disentangled representation of the data and having humans decide which features should not have been used for the task.
-For a meaningful comparison, we show in Tab.
-1 existing bias identification methods, highlighting aspects we consider fundamental for the bias identification task.
-Debiasing Debiasing and bias extraction techniques have become crucial in ensuring the fairness and accuracy of machine learning models [36], with extensive research dedicated to removing harmful biases across various domains.
-Some existing methods use bias annotations to train unbiased models, by means of group balanced subsampling [14], reweighting [33] or data augmentations [44].
-In the absence of these annotations, other works [25, 20, 28] have proposed to first learn a biased model and then focus on its mistakes to train an unbiased one.
-Fairness Fairness in has been extensively studied, with numerous approaches [8, 38] proposed to facilitate ethical research and to ensure equitable outcomes across different subpopulations.
-Most of those methods overlap with domain generalization and worst-group performance improvements.
-This is also a field where model interpretability plays a crucial role [32], as understanding how decisions are made can help in identifying and mitigating biases.
-Invariant Learning Robustness to out-of-distribution changes can be obtained by enforcing that the learning model is invariant to different environments or domains [2, 16, 44].
-But there are many cases where we don’t have access to such labeled environments and we must discover them.
-Approaches like [9] partition the data into subsets that maximally contradict an invariant constraint, and apply algorithms for distributional robustness, like groupDRO [33], on those subsets, called environments.
-Model guidance with explanations Most similar to our bias-based regularization is the use of model explanations to guide learning [31, 35, 27, 42].
-Using differentiable attribution methods like Grad-CAM [34], they constrain the attention of the trained models to only fall on the class instances, segmented at the input level.
-A limitation of these approaches is that they can’t prevent the model from learning biases that are not spatially separated from the class in the input (e.g.
-color, shape, or accessories of the subject).
+    We show that our method is able to expose previously untapped ImageNet-1k spurious correlations, and further proceed to show that multiple state-of-the-art models are affected by them, highlighting their inability to formulate robust definitions for the targeted classes.
+2 Background
+Machine learning methods naturally capture relevant factors needed to solve a task.
+However, models might also capture shortcuts (Geirhos et al., 2020), as correlations between non-essential features of the inputs and the label.
+These shortcuts represent spurious correlations, that don’t hold in a more general setup (e.g.
+using water features to classify waterbirds instead of focusing on the birds’ features), and should not be used for reliable generalization outside the training distribution, as they often lead to degraded performance (Quiñonero-Candela et al., 2008; Beery et al., 2018; Hendrycks et al., 2021).
+SCs from error analysis
+Approaches like B2T (Kim et al., 2024) rely exclusively on the validation samples, identifying which correlations between concepts are more prevalent in the misclassified examples.
+To catch SCs, B2T needs samples that oppose the strong correlations in the training set, thus leading to misclassification.
+To circumvent this limitation, DrML (Zhang et al., 2023b) manually builds a list of texts containing classes and concepts associations, that could potentially underline an SC.
+It forwards each such textual association through a classifier, learned on the image modality, and keeps as SCs the erroneously classified ones.
+In those approaches, the burden falls on the practitioners to come up with exhaustive samples or associations, failing to detect unexpected SCs.
+Differently, WASP focuses on analyzing the explicit learned weights of a model, covering all trainset samples.
+We thus extract the spuriously correlated concepts directly from the weights, bypassing the need for an exhaustive validation set or correlations candidates.
+SCs from train data analysis
+In SpLiCE (Bhalla et al., 2024) each image is decomposed into high-level textual concepts, searching next for concepts that are frequent for a certain class, but not for the others.
+LG (Zhao et al., 2024) relies on LLMs to propose concepts potentially correlated with each class, using image captions.
+Next, it uses CLIP (Radford et al., 2021) to estimate a class-specificity score for each concept, and highly scored concepts for a class w.r.t.
+the others are considered SCs.
+These methods focus on the concepts’ occurrences per class, making them prone to missing low-frequency concepts, as their presence can be drowned when averaging scores over a large dataset.
+Moreover, the SCs found through data analysis could be harder to learn than the class itself, so they are not necessarily imprinted in the model.
+In contrast, learned SCs (including error analysis revealed ones) must always be addressed, as they are, by definition, proven to impact the classifier.
+For this reason, WASP targets learned SCs by looking directly at the impact of the training set upon the model’s weights.
+Manual interpretation of correlations
+The method introduced in (Singla & Feizi, 2021) finds spurious features learned by a model, but it requires humans to manually annotate whether an image region is causal or not for a class.
+While this ensures a higher quality of the annotations, it also poses problems of scalability to large datasets.
+In contrast, WASP works fully automated, at scale, identifying SCs for each class in ImageNet-1k.
+SCs from subpopulation shift setup Other previous works (Pezeshki et al., 2024; Liu et al., 2021a; Arefin et al., 2024; Zare & Nguyen, 2024) have focused on SC identification strictly within the context of subpopulation shifts.
+The particularity of this setup is that the training and validation sets always contain subsets of samples that oppose the strong spurious correlations of the dataset.
+With WASP, we do not have this limitation.
+Most of these methods (Pezeshki et al., 2024; Liu et al., 2021a; Zare & Nguyen, 2024) focus on first learning a strongly biased classifier and then either separate the samples of each class into two groups (Pezeshki et al., 2024; Zare & Nguyen, 2024) (one containing correctly classified examples and the other containing misclassified ones), or place higher weights on hard samples (Liu et al., 2021a), in order to balance the dataset.
+CoBalT (Arefin et al., 2024) on the other hand uses an unsupervised method for object recognition and then samples the dataset examples such that all object types are uniformly distributed in each class.
+Their result contains heatmaps overlays on images, which can offer insights to guide further manual SCs identification.
+Fairness It is important to note that the proposed method can be utilized to evaluate the fairness of a given dataset and that we do conduct benchmarking on the CivilComments dataset, which encompasses racial and religious concerns.
+However, it is critical to emphasize that our approach is neither designed to measure nor address issues of fairness.
+Instead, our method is specifically developed to examine whether a given dataset imparts a clear definition of the featured classes to a model — namely, whether classifiers learn spurious correlations and confound class features with environmental features.
+Accordingly, our work is situated within the literature on subpopulation shift setups and we assess the quality of our proposed approach within this framework.
+Evaluating our approach on fairness benchmarks lies outside the scope of the current study, but may constitute a subject for subsequent research.
 3 Our Method
-For a standard classification task, we aim to identify biases acquired by a model through training on a new dataset.
-By dataset biases we refer to class-independent concepts, whose presence in the samples greatly affects the class label distribution.
+For a standard classification task, we aim to identify spurious correlations learned by a model through training on a new dataset.
+By dataset’s SCs we refer to class-independent concepts, whose presence in the samples greatly affects the class label distribution.
 By concepts we refer to words or expressions with a semantic load, that are most relevant in describing a given context.
-For instance, to extract concepts from an image dataset, we begin with the image captions, extract keywords from them, and use the keywords as concepts.
-Through biases of a classifier $f_{\theta}$ we refer to concepts causally unrelated with a class, whose presence in the input significantly changes the distribution of class probabilities predicted by $f_{\theta}$.
-We further classify them as positively correlated biases w.r.t.
-to a class $k$ if the presence of these biases in the input increases the probability of $f_{\theta}$ predicting the class $k$ and negatively correlated biases if they decrease it.
-Setup We start with a dataset with samples $(x_{j},y_{j})$, in which we search for biases, through the representations offered by a foundation model $M$, capable of embedding both the input samples $x_{j}$ and the concepts $c_{i}$ into the $\mathbb{R}^{D}$ vector space.
+Through SCs learned by a classifier $f_{\theta}$ we refer to concepts causally unrelated with a class, whose presence in the input significantly changes the distribution of class probabilities predicted by $f_{\theta}$.
+We further classify them as positively correlated concepts w.r.t.
+to a class $k$, if their presence in the input increases the probability of $f_{\theta}$ predicting the class $k$, and negatively correlated concepts if they decrease it.
+Setup We start with a dataset with samples $(x_{j},y_{j})$, in which we search for SCs, through the representations offered by a foundation model $M$, capable of embedding both the input samples $x_{j}$ and the concepts $c_{i}$ into the $\mathbb{R}^{D}$ vector space.
 We further assume that their cosine similarity $cos(M(c_{i}),M(x_{j}))$ is higher when the concept $c_{i}$ is present in sample $x_{j}$ than when it is not.
 We first construct a set of concepts $(c_{i})_{1\leq i\leq q}$ in textual form, that are present in the training data.
-To spot the biases among them, we train a linear layer on top of the input data representations given by $M$.
-Next, through concept filtering, we identify concepts that are causally unrelated to the classes, revealing our bias candidates.
-Finally, we rank them based on how strongly they influence the predictions of this linear classifier and automatically decide the bias score cut-off.
+To spot the spuriously correlated ones among them, we train a linear layer on top of the input data representations given by $M$.
+Next, through concept filtering, we identify concepts that are causally unrelated to the classes, revealing our candidates.
+Finally, we rank them based on how strongly they influence the predictions of this linear classifier and automatically decide the SC score cut-off.
 In the following equations, we use the dot product instead of the cosine similarity, as all the vectors we use are normalized.
 The main steps of our method (also revealed in Fig.
-1, and in more detail in the Appx.
-C algorithm), are the following:
+2, and in more detail in the Appx.
+D algorithm), are the following:
 Initialization We initialize $w_{k}$, the weights of class $k$ in the linear layer, with the embedding of its corresponding class name, extracted by the model $M$:
 $w^{0}_{k}=M(\text{class\_name}_{k});\;\forall 1\leq k\leq|\mathcal{Y}|=N.$ (1)
-Figure 2: The maximum distance between the reference line and the smoothed bias scores gives the threshold for our dynamic cut-off heuristic.
-Step 1: Model training We train the weights using ERM [37] on our dataset of interest.
+Figure 3: The maximum distance between the reference line and the smoothed scores gives the threshold for our cut-off heuristic.
+Step 1: Model training We train the weights using ERM (Vapnik, 1999) on our dataset of interest.
 Through learning, the class concepts (represented by the weights in the linear layer) naturally shift from the original initialization towards concepts that can effectively distinguish between class samples, given the current dataset, as visually presented in Fig.
-1.
-This includes the classes themselves and their (positive) biases as well.
-At the end of the training, the weights can be expressed as a weighted sum of these concepts.
-Step 2: Bias identification Using the learned weights, we continue to identify which concepts $c_{i}$ are either biases positively or negatively correlated with each class $k$.
+2.
+This includes not only the classes themselves but also concepts that are positively correlated with them.
+At the end of the training, the weights can be expressed as a weighted sum of classes and concepts.
+Step 2: SCs identification Using the learned weights, we continue to identify which concepts $c_{i}$ are either positively or negatively correlated with each class $k$.
 First, we filter out the class-related concepts.
-We then rank the remaining class-neutral concepts and select the top ones as biases based on a dynamic thresholding strategy, as follows:
+We then rank the remaining class-neutral concepts and mark the top ones as SCs, based on a dynamic thresholding strategy, as follows:
 a.
-Filter concepts We first aim to filter out concepts that are indeed related to a certain class, leaving only concepts that are causally unrelated to a class (we call them class-neutral concepts).
-These concepts are proper candidates for biases because they are not necessary for the recognition of a class (e.g.
+Filter out class-related concepts Initially, we extract dataset’s concepts using existing tools.
+We first aim to filter out concepts that are indeed related to a certain class, leaving only concepts that are causally unrelated to a class (we call them class-neutral concepts).
+These concepts are proper candidates for SCs because they are not required for the recognition of a class (e.g.
 a forest background is well correlated with species of landbirds, but we want to prevent the model from using this causally unrelated correlation).
@@
 4.
-b. Rank class-neutral concepts For each class-neural concept $c_{i}$ and class $k$, we rank the concepts based on their similarity, using the following positive-bias score:
+b. Rank class-neutral concepts For each class-neural concept $c_{i}$ and class $k$, we rank the concepts based on their similarity, using the following positive-SC score:
 $s_{k,i}^{+}=w^{\top}_{k}M(c_{i})-{\underset{k^{\prime}=\overline{1,N}}{\text{min}}}w^{\top}_{k^{\prime}}M(c_{i}).$ (2)
 Intuitively, we want concepts specific to one class but not to all the others.
-For this, for each class $k$, we select concepts highly correlated to it, while not correlated to at least one other class.
-For the negatively correlated biases, the score $s_{k,i}^{-}$ is similar, with the difference that we are taking into account the dissimilarity scores ($-w^{\top}_{k}M(c_{i})$).
-Table 2: Validation procedure for three positively correlated biases found using ConceptDrift for CLIP ViT-L/14 fine-tuned on ImageNet.
-We evaluate the model’s capability to recognize a depicted (correct) class before and after the introduction of a biasing factor.
+For this, for each class $k$, we select the concepts which starkly correlate to it, when compared to other classes.
+For the negatively correlated SCs, the score $s_{k,i}^{-}$ is similar, with the difference that we are taking into account the dissimilarity scores ($-w^{\top}_{k}M(c_{i})$).
+Table 1: Validation procedure for three positively correlated concepts found using WASP for CLIP ViT-L/14 fine-tuned on ImageNet.
+We evaluate the model’s capability to recognize a depicted (correct) class before and after the introduction of an identified concept.
 For each prompt, $1000$ images are generated using FLUX.1-dev.
-We observe throughout all considered scenarios, a significant drop in the model’s capacity to identify the correct class when the biasing factor is involved.
+We observe throughout all considered scenarios, a significant drop in the model’s capacity to identify the correct class when the selected concept is involved.
 We also observe a meaningful increase in the likelihood of having the induced class predicted, even though it is not illustrated.
-Exploited Bias $\rightarrow$ Avg. Softmax Activations (%) Samples Predicted As (%)
-Induced Class Correct Class Prompt Correct Class Induced Class Correct Class Induced Class
-firemen (bias) $\rightarrow$ fire truck (class) peafowl a photo of a peafowl 99.9 0.0 100.0 0.0
-firemen and a peafowl 10.6 (-89.3 ) 78.9 (+78.9) 5.3 (-94.7) 93.4 (+93.4)
-reading a newspaper (bias) $\rightarrow$ crossword (class) Mexican hairless dog a photo of a mexican hairless dog 45.2 0.0 47.5 0.0
-a man reading a newspaper in a chair with a mexican hairless dog in its lap 2.0 (-43.2) 21.6 (+21.6) 0.9 (-46.6) 36.6 (+36.6)
-shrimp (bias) $\rightarrow$ American lobster (class) Bernese Mountain Dog a photo of a Bernese Mountain Dog 90.2 0.0 99.8 0.0
-shrimp and pasta near a Bernese Mountain Dog 8.1 (-82.1) 14.8 (+14.6) 10.6 (-89.2) 37.2 (+37.2)
+Correct Exploited SC Prompt Avg. Softmax Activations (%) Samples Predicted As (%)
+Class (Induced Class) Correct Class Induced Class Correct Class Induced Class
+peafowl firemen $\bullet$ a photo of a peafowl 99.9 0.0 100.0 0.0
+(fire truck) $\bullet$ firemen and a peafowl 10.6 (-89.3 ) 78.9 (+78.9) 5.3 (-94.7) 93.4 (+93.4)
+Mexican reading a $\bullet$ a photo of a Mexican hairless dog 45.2 0.0 47.5 0.0
+hairless newspaper $\bullet$ a man reading a newspaper in a chair 2.0 (-43.2) 21.6 (+21.6) 0.9 (-46.6) 36.6 (+36.6)
+dog (crossword) with a Mexican hairless dog in his lap
+Bernese shrimp $\bullet$ a photo of a Bernese Mountain Dog 90.2 0.0 99.8 0.0
+Mountain (American $\bullet$ shrimp and pasta near a Bernese 8.1 (-82.1) 14.8 (+14.8) 10.6 (-89.2) 37.2 (+37.2)
+Dog lobster) Mountain Dog
+Table 2: Accuracy of various convolutional and transformer-based models trained on ImageNet-1k, on the data generated for Tab.
+1.
+As with Fig.
+1 and Fig.
+4, we note that the performance of these models is significantly affected, even though the correct class is illustrated right in front and center while the predicted class is absent from the generated images.
+An exhaustive list is presented in Appx.
+B.
+Prompt employed (correct class highlighted in bold, SC in red)
+Model a photo of a peafowl firemen and a peafowl a photo of a Bernese Mountain Dog shrimp and pasta near a Bernese Mountain Dog
+alexnet (Krizhevsky, 2014) $100.0$ $4.6$ (-95.4) $96.2$ $23.3$ (-72.9)
+efficientnet_b1 (Tan & Le, 2019) $100.0$ $42.6$ (-57.4) $88.1$ $67.1$ (-21.0)
+regnet_x_32gf (Radosavovic et al., 2020) $100.0$ $66.1$ (-33.9) $85.9$ $46.0$ (-39.9)
+resnet50 (He et al., 2016) $100.0$ $30.1$ (-69.9) $73.9$ $54.5$ (-19.4)
+resnext101_32x8d (Xie et al., 2017) $100.0$ $66.6$ (-33.4) $84.7$ $61.2$ (-23.5)
+squeezenet1_1 (Iandola et al., 2016) $100.0$ $13.8$ (-86.2) $91.2$ $46.1$ (-45.1)
+swin_b (Liu et al., 2021b) $100.0$ $81.5$ (-18.5) $95.2$ $72.6$ (-22.6)
+vgg19_bn (Simonyan & Zisserman, 2015) $100.0$ $35.9$ (-64.1) $83.1$ $46.2$ (-36.9)
+vit_l_16 (Dosovitskiy et al., 2021) $100.0$ $55.9$ (-44.1) $95.3$ $76.0$ (-19.3)
+wide_resnet50_2 (Zagoruyko & Komodakis, 2016) $100.0$ $60.6$ (-39.4) $95.7$ $63.9$ (-31.8)
 c.
-Dynamic thresholding towards biases We select the biases of each class using a dynamic threshold for the bias scores above.
-We keep the highest-ranked ones, both for positive and negative bias scores.
+Dynamic thresholding We select the SCs of each class using a dynamic threshold for the scores above, keeping the highest-ranked ones.
 We remove the null scores and then apply a mean filter with a window size of $r$ to smooth the curve of scores.
 We denote the scores obtained at this step with $(\overline{s}_{k,i})_{1\leq i\leq p}$, with $p=q-r+1$.
-We then select the top $m$ ones, as positive biases for class $k$, where the $m$ index is defined as:
+We then select the top $m$ ones, as positive SCs for class $k$, where the $m$ index is defined as:
 $m=\lfloor r/2\rfloor+\arg\max_{i}\left(\overline{s}_{k,1}-i\frac{\overline{s}_{k,1}-\overline{s}_{k,p}}{p-1}-\overline{s}_{k,i}\right).$ (3)
-The intuition is that $m$ represents the index where the smoothed score curve deviates most from the straight reversed diagonal line, connecting $\overline{s}_{k,1}$ and $\overline{s}_{k,p}$ (see Fig 2).
-Figure 3: Qualitative results depicting a series of biases uncovered employing our proposed ConceptDrift approach for CLIP ViT-L/14 fine-tuned on ImageNet [10].
+The intuition is that $m$ represents the index where the smoothed score curve deviates most from the straight reversed diagonal line, connecting $\overline{s}_{k,1}$ and $\overline{s}_{k,p}$ (see Fig.
+3).
+Figure 4: Qualitative results depicting a series of SCs, uncovered employing our proposed WASP approach for CLIP ViT-L/14 fine-tuned on ImageNet (Deng et al., 2009).
 All images are generated using FLUX.1-dev.
-In each scenario, a (correct) ImageNet class is depicted alongside a bias found to be positively correlated with a different (induced) class.
-Alongside each image, we provide the names of the correct and predicted classes, together with the softmax activations of their respective neurons.
-Though in most scenarios a single ImageNet class is clearly depicted, the model chooses to ignore it and predict a class that is not illustrated, clinging to the spurious correlations learned.
+In each scenario, a (correct) ImageNet class is depicted alongside a concept found to be positively correlated with a different (induced) class.
+We provide the names of the correct and predicted classes (which are specifically those induced by the chosen spuriously correlated concept).
+Though in most scenarios a single ImageNet-1k class is clearly depicted, the model ignores it and predicts a class that is not illustrated, clinging to the spurious correlations learned.
 4 Experimental Setup
-Foundation models (FM) We used mGTE (gte-large-en-v1.5 [45]) for text embeddings in CivilComments [5], and OpenAI CLIP ViT-L/14 [30] for text and images otherwise.
-Concept extraction For the image classification task, we first use the GIT-Large  [41] captioning model (trained on MSCOCO [19]) to obtain descriptions of the dataset’s images.
-Next, we extract concepts from the captions (or directly from the text samples for the text classification task), using YAKE [7] keyword extractor, taking the top $256$ n-grams for $n=3,5$.
-Concepts filtering We use Llama-3.1-8B-Instruct [11] to remove class instances from the concepts extracted at the previous step.
-We also apply post-processing using WordNet [24] to catch obvious class instances that the LLM might miss.
-To do this, for each class we specify a name used to search for synsets in the WordNet [24] hierarchy (e.g.
-bird for Waterbirds) and then remove individual words that match with any hyponym or hypernym of those synsets.
-Training We train the linear layer on $L_{2}$ normalized embeddings extracted by these models using the PyTorch [26] AdamW [23] optimizer with a learning rate of $41e-$ 4, a weight decay of $51e-$ 5 and a batch size of $1024$.
+Foundation models (FM) We used mGTE (gte-large-en-v1.5 (Zhang et al., 2024)) for text embeddings in CivilComments (Borkan et al., 2019), and OpenAI CLIP ViT-L/14 (Radford et al., 2021) for text and images otherwise.
+Concept extraction For the image classification task, we first use the GIT-Large  (Wang et al., 2022) captioning model (trained on MSCOCO (Lin et al., 2014)) to obtain descriptions of the dataset’s images.
+Next, we extract concepts from the captions (or directly from the text samples for the text classification task), using YAKE (Campos et al., 2020) keyword extractor, taking the top $256$ n-grams for $n=3,5$.
+Concepts filtering We use Llama-3.1-8B-Instruct (Dubey et al., 2024) to remove class instances from the concepts extracted at the previous step.
+We also apply post-processing using WordNet (Miller, 1995) to catch obvious class instances that the LLM might miss.
+To do this, for each class we specify a name used to search for synsets in the WordNet (Miller, 1995) hierarchy (e.gbird for Waterbirds) and then remove individual words that match with any hyponym or hypernym of those synsets.
+Training We train the linear layer on $L_{2}$ normalized embeddings extracted by these models using the PyTorch (Paszke et al., 2017) AdamW (Loshchilov & Hutter, 2019) optimizer with a learning rate of $41e-$ 4, a weight decay of $51e-$ 5 and a batch size of $1024$.
 We use the cross entropy loss with balanced class weights as the objective.
-The weights of the layer are normalized after each update and we use CLIP’s [30] temperature to scale the logits.
+The weights of the layer are normalized after each update and we use CLIP’s (Radford et al., 2021) temperature to scale the logits.
 We use the class-balanced accuracy on the validation set as the criterion for model selection and early stopping.
 4.1 Datasets
-Waterbirds [40] is a common datasets for generalization and bias mitigation.
-It is created from CUB [39], by grouping species of birds into two categories, landbirds and waterbirds, each being associated with a spurious correlation regarding its background, land and water respectively.
-CelebA [21] is a large-scale collection of celebrity images (over 200000), widely used in computer vision research.
-The setup for using it in a generalization context [22] consists of using the Blond_Hair attribute as the class label and the gender as the spurious feature.
-CivilComments [5] is a large collection comprised of $1.8$ million online user comments.
+Waterbirds (Wah et al., 2011b) is a common dataset for generalization and mitigating spurious correlations.
+It is created from CUB (Wah et al., 2011a), by grouping species of birds into two categories, landbirds and waterbirds, each being associated with a spurious correlation regarding its background, land, and water respectively.
+CelebA (Liu et al., 2015a) is a large-scale collection of celebrity images (over $200.000$), widely used in computer vision research.
+For generalization context, the setup (Liu et al., 2015b) consists of using the Blond_Hair attribute as the class label and the gender as the spurious feature.
+CivilComments (Borkan et al., 2019) is a large collection comprised of $1.8$ million online user comments.
 This dataset is used employed in NLP bias and fairness research concerning different social and ethnical groups.
-ImageNet [10] is a larger-scale popular dataset for image classification ($1000$ classes, with approximately $1300$ training samples and $50$ validation samples per class).
-5 Bias identification evaluation
-For a proper quantitative evaluation of our proposed biases, beyond the subjectivity of the qualitative aspects, we combine the bias identified by ConceptDrift with different components: in Sec.
-5.1 we generate samples exploiting the discovered biases; in Sec.
-5.2 we assess GroupDRO [33] robustness over training set partitions based on the proposed biases groups; in Sec.
-5.3 we evaluate bias-based training regularization; and in Sec.
-5.4 we use a minimally bias-enhanced prompting for zero-shot prediction with a Foundational Model.
-An exhaustive list of the extracted biases for all datasets can be found in Appx.
-B.
-5.1 ImageNet biases
-We employ ConceptDrift to uncover a series of biases and spurious correlations plaguing the decision-making process of OpenAI’s CLIP ViT-L/14 fine-tuned on ImageNet.
-As the biases our method reveals lie beyond the restricted scope of the validation set, our sole recourse for illustrating and validating these findings at scale lies in generating synthetic images.
-To this end, we have employed a quantized version of FLUX.1-dev [17].
-To validate the impact of the biases, we prompt the generative model to depict a chosen (correct) class alongside a non-ImageNet object that we identified as a positively correlated bias for a different class.
-The validation process is presented for three distinct scenarios in Tab.
-2.
-Each scenario is defined by a pair of classes: a correct class which is illustrated in the image and another class which is absent, but induced through the presence of a biasing object or activity.
-We measure the impact of the bias by means of comparing the model’s ability to predict the correct class before and after the introduction of the biasing factor.
-We generate $1000$ images for each prompt, and we evaluate the model both in terms of the predicted class and in terms of the activations of the class neurons.
-We observe throughout all considered scenarios, a significant drop in the model’s capacity to identify the correct class when the biasing factor is involved, and a significantly increased likelihood to have the induced class predicted, even though it is not illustrated in any way, shape or form in the generated image.
-Within the same context, we present a series of qualitative examples in Fig.
-3.
-We emphasize that, even though throughout most of these samples, a single ImageNet class is clearly depicted, the model chooses to ignore it and label the image as a completely different class, not illustrated at all in the image, solely based on the presence of a non-ImageNet object.
-As an example, the network chooses to completely ignore the peafowl, present right at the center of the second image, and labels the image as a ”Fire Truck” based on the presence of the firefighters.
-The model at hand is generally considered to be a robust state-of-the-art model, benefiting from ample pre-training, achieving an $82.4\%$ top-1 accuracy on the ImageNet validation set.
-We emphasize through our experiment that even under these reassuring circumstances, critical reasoning flaws can make their way through in a production-ready model, undetected by validating and examining the model’s performance on held-out data (the traditional approach for identifying biases).
-It is through shifting the approach towards investigating the model’s weights that we can anticipate and prevent serious biases from making their way into the final version of the model.
-Throughout this validation process, we have manually ensured that the generated data does not contain troubling artifacts.
-For instance, it is the case that even though the prompt only asks for a firefighter, in certain circumstances the generative network will also depict a fire truck alongside the other features requested by the prompt, against our desires.
-Additionally, sometimes the generative model is not precise and may depict a dog or a bird species different than the one we asked for.
-These examples of faulty generation are not exhaustive.
-It is therefore the case that prompts have to be carefully designed and images carefully inspected, in order to ensure the quality of the evaluation.
-Under these circumstances, we have endeavored to illustrate and validate but a small set of the remarkable flaws, ascribing a complete examination of our findings to future work.
-However, the complete list of ranked concepts for each class, determined automatically by employing our ConceptDrift approach, is available for consultation in Appx.
-B.
-Table 3: Learning in the context of 100% spurious correlations.
-In the absence of samples with negatively correlated biases, GroupDRO performs similarly with the standard ERM.
-In contrast, bias regularization using our biases yields excellent outcomes in this setup, improving the worst group accuracy above the considered baselines: ERM, GroupDRO and the bias regularization with random causally-unrelated concepts (obtained after filtering in Step2a.).
+ImageNet (Deng et al., 2009) is a larger-scale popular dataset for image classification ($1000$ classes, with approx. $1300$ training samples and $50$ validation samples per class).
+5 Evaluating spurious correlations
+For a proper quantitative evaluation of our proposed SCs, beyond the subjectivity of the qualitative aspects, we combine the concepts identified by WASP with different components: in Sec.
+5.1 we generate samples exploiting the discovered concepts; in Sec.
+5.2 we evaluate in scenarios lacking spurious correlation counterexamples; and in Sec.
+5.3 we use a minimally SC-enhanced prompting for zero-shot prediction with a foundation model.
+A list of the extracted concepts for all datasets can be found in Appx.
+C.
+Table 3: Learning in the context of perfect spurious correlations.
+In the absence of samples that associate a class instance and concepts spuriously correlated with other classes, GroupDRO does not outperform the standard ERM.
+In contrast, our regularization based on the identified concepts consistently yields improvements (concerning worst group accuracy) over the considered baselines: ERM, GroupDRO, and the regularization with random causally unrelated concepts (obtained after the filtering in Step2a).
 Waterbirds (Acc % $\uparrow$) CelebA (Acc % $\uparrow$) CivilComments (Acc % $\uparrow$)
@@
 GroupDRO 38.9$\pm$5.4 71.2$\pm$2.0 8.1$\pm$0.3 60.3$\pm$1.0 18.7$\pm$0.4 50.2$\pm$0.5
-Bias regularization w/ random concepts (class-neutral) 46.6$\pm$2.7 75.3$\pm$1.1 9.4$\pm$0.0 61.4$\pm$2.0 19.1$\pm$1.6 50.8$\pm$0.9
-Bias regularization w/ ConceptDrift biases (ours) 57.9$\pm$0.3 79.8$\pm$0.1 10.4$\pm$0.5 62.0$\pm$1.8 31.3$\pm$0.7 57.5$\pm$0.4
-5.2 From biases to groups, towards generalization
-Next, we move to the subpopulation shift setup [33] (where different biases determine subgroups in the data) and follow the evaluation protocol in SubpopBench [43].
-Normally, ERM training on such data behaves poorly on minority groups, but methods like GroupDRO [33] can recover the performance, when the subgroups are given.
-We show that the biases discovered by our method can be used to create such subgroups, proving that they are indeed spuriously correlated concepts that a robust model should not rely on.
-Table 4: Generalization task, evaluated by comparing the performance of an ImageNet-pretrained ResNet-50 model, trained with GroupDRO on dataset partitions that are based on the identified biases.
-ConceptDrift biases are competitive with SoTA.
-.
-Waterbirds (Acc % $\uparrow$) CelebA (Acc % $\uparrow$) Training Bias identification Worst Avg.
-Worst Avg.
-ERM - 70.0$\pm$2.4 84.9$\pm$0.9 54.2$\pm$3.6 95.4$\pm$0.3 GroupDRO w/ B2T [15] partitioning into groups B2T [15] 83.2±$\pm 3.6$ 90.6±$\pm 2.0$ 87.6±$\pm 2.0$ 93.7±$\pm 0.4$ Lg [47] 81.4±$\pm 3.4$ 89.5±$\pm 1.3$ 89.3±$\pm 1.8$ 93.2±$\pm 0.6$ ConceptDrift (ours) 81.9±$\pm 2.2$ 90.0±$\pm 2.3$ 88.4±$\pm 1.4$ 93.6±$\pm 0.4$ Ground-truth groups 85.8±$\pm 2.5$ 90.1±$\pm 1.6$ 88.8±$\pm 2.0$ 93.4±$\pm 0.7$
-For the experiments in Tab.
-4, we use the following pipeline.
-Based on the identified biases, we split the training data into subgroups with the partitioning technique from B2T [15].
-Next, we train the model with GroupDRO, constraining it to attend equitably to the subgroups.
-If the concepts represent biases that will lead the model to learn spurious correlations through ERM, then under GroupDRO training, the model will reduce its reliance on them, surpassing the performance of ERM on the minority subgroups.
-Tab.
-4 shows that our identified ConceptDrift-biases are indeed useful for GroupDRO, far surpassing the ERM baseline and obtaining results that are competitive with the other bias identification methods in terms of worst-group performance on the Waterbirds and CelebA datasets.
-Note that the partitioning method was not tuned to accommodate our biases (e.g.
-our biases include actions and attributes, while the other methods only propose nouns).
-Loss correlates with the presence of biases In the setup described above, we look at the correlations between the loss values and the bias-to-sample similarities.
-We compare ERM with GroupDRO applied on subgroups obtained with our biases and the B2T partitioning strategy.
-See in Appx.
-D how after GroupDRO with our proposed biases, the correlations significantly decrease, revealing that the model is less prone to make mistakes on the biased samples.
-Intuitively, this also indicates that the model is less reliant on the biases.
-5.3 Biases used for training regularization
-We previously showed in Sec.
-5.2 that partitioning datasets using qualitative biases, followed by GroupDRO training, can significantly improve the robustness of a model to the bias-induced distribution shifts.
-We now investigate a scenario where prior methods for robustness like GropuDRO or others [33, 44, 14] are not applicable.
-In a setup with 100% bias-to-class correlations, the absence of the diverse bias-class associations renders these existing methods ineffective.
-We thus propose the task of learning in this perfect correlated setup, removing all bias conflicting samples from the training and validation sets.
-We use the discovered biases to impose regularization upon the trained models, in particular, regularization for linear probing, learned on top of frozen CLIP embeddings.
-Intuitively, we constraint each bias to be equally distanced from the class weights.
-We formulate this as a MSE between the similarity of $b$ with a class weight and the average similarity of $b$ to all the class weights, scaled by $\tau$ (CLIP’s temperature):
+Regularization w/ random concepts (class-neutral) 46.6$\pm$2.7 75.3$\pm$1.1 9.4$\pm$0.0 61.4$\pm$2.0 19.1$\pm$1.6 50.8$\pm$0.9
+Regularization w/ SCs found by Lg 50.4$\pm$0.1 76.6$\pm$0.0 8.3$\pm$0.0 61.2$\pm$0.5 - -
+Regularization w/ SCs found by WASP (ours) 57.9$\pm$0.3 79.8$\pm$0.1 10.4$\pm$0.5 62.0$\pm$1.8 31.3$\pm$0.7 57.5$\pm$0.4
+5.1 ImageNet Spurious Correlations
+We employ WASP to uncover a series of spurious correlations plaguing the decision-making process of OpenAI’s CLIP ViT-L/14 fine-tuned on ImageNet.
+As the SCs our method reveals lie beyond the restricted scope of the validation set, our sole recourse for illustrating and validating these findings at scale lies in generating synthetic images.
+To this end, we have employed a quantized version of FLUX.1-dev (Labs, 2024).
+To validate the impact of the SCs, we prompt the generative model to depict a chosen (correct) class alongside a non-ImageNet object that we identified as a positive SC for a different class.
+The validation process is presented for three distinct scenarios in Tab.
+1.
+Each scenario is defined by a correct class that is illustrated in the image, a concept (object, property, or activity) that is not causally tied to any class, and an absent class induced through the presence of the concept.
+We expect the classifier to predict this absent class, based on our scoring.
+We measure the impact of the concept by comparing the model’s ability to predict the correct class before and after its introduction.
+We generate $1000$ images for each scenario and we evaluate the model both in terms of accuracy and in terms of the frequency with which it predicts the induced class.
+Throughout all considered scenarios, we observe a significant drop in the model’s capacity to identify the correct class when the biasing factor is involved, and a significantly increased likelihood of having the induced class predicted, even though it is not illustrated in any way, shape or form in the generated image.
+Within the same context, we present a series of qualitative examples in Fig.
+4.
+We emphasize that, even though throughout most of these samples, a single ImageNet-1k class is clearly depicted, the model chooses to ignore it and label the image as a completely different class, not illustrated at all in the image, solely based on the presence of a non-ImageNet object.
+We underline, by means of the results presented in Tab.
+1, that the model is not fooled by artifacts in the generated images to predict randomly.
+We test the performance of the models on images featuring the correct class, without added objects.
+We observe this way that the model’s performance on the generated data is on par with the original performance of the model on these classes, validating that the generated images are not out of distribution.
+Furthermore, we show that the rate at which the induced class is predicted increases significantly, but more importantly, that the presence of the SC significantly increases the softmax activations of the model for the induced class (considering that it is spread across 1000 classes).
+The model at hand is generally considered to be a robust state-of-the-art model, benefiting from ample pre-training, achieving an $82.4\%$ top-1 accuracy on the ImageNet validation set.
+We emphasize through our experiment that even under these reassuring circumstances, critical reasoning flaws can make their way through, in a production-ready model, undetected by validating and examining the model’s performance on held-out data.
+We further proceed to examine the impact of the SCs for which we have generated data on an exhaustive set of ImageNet-1k state-of-the-art models.
+We present the entire set of results in Appx.
+B, and the results for a selection of these models in Tab.
+2.
+Throughout this validation process we had to manually ensure that the data is generated correctly.
+Thus, we were forced to restrict ourselves to presenting only a few of our findings, ascribing a complete investigation of the impact of all SCs found to future work.
+We present a set of tables featuring the scores computed for different concepts with respect to different classes in Appx.
+C.
+5.2 Training in a Fully Spurious Setup
+We explore here an extreme setup, featuring no spurious correlation counterexample.
+In real world, this might be the case for most of the SCs, since usually this kind of correlations are generated by decisions in dataset acquisitions and filtering.
+Removing the minority groups from common robustness datasets, renders GroupDRO-like approaches completely ineffective, as their performance at best only matches the standard ERM (see Tab.
+3).
+We use the employed SCs to impose a regularization upon the trained linear probes, learned on top of frozen embeddings from the FM.
+Intuitively, we constraint the weights to be equally distanced from the identified SCs.
+We formulate this as an MSE between the similarity of class weight $w_{k}$ with an SC $b$ and the average similarity of all class weights with $b$, these terms being scaled by $\tau$ (CLIP’s temperature):
 $\mathcal{L}_{reg}(b)=\frac{\tau^{2}}{N}\sum^{N}_{k=1}\left[w^{\top}_{k}M(b)-sg\left(\frac{1}{N}\sum^{N}_{j=1}w^{\top}_{j}M(b)\right)\right]^{2},$ (4)
@@
 $\mathcal{L}=\mathcal{L}_{ERM}+\alpha\frac{1}{\left|\mathcal{B}\right|}\sum_{b\in\mathcal{B}}\mathcal{L}_{reg}(b),$ (5)
-where $\mathcal{B}$ is the set of selected biases and $\alpha=0.1$.
+where $\mathcal{B}$ is the set of selected concepts and $\alpha=0.1$.
 In Tab.
-3, we present the results of linear probing with $\mathcal{L}$ loss, in this perfect spurious correlation scenario.
-The improvement in worst group accuracy shows that the learned classification weights are less reliant on the biases, thus more robust and better aligned with the classification task.
-Table 5: Results for the Foundation Model (FM) Zero-shot prompting task.
-We help the model account for biases by explicitly introducing them in the prompting scheme.
-We perform tests employing biases discovered by different methods and evaluate the zero-shot performance of the FM.
-Aided by the biases discovered through our proposed method, the model’s performance, measured in terms of worst group accuracy, is significantly enhanced across all test scenarios.
-Our method surpasses in this regard all previous bias-identification approaches.
+3, we present the results of linear probing with this loss, in the previously mentioned scenario, with no SC counterexamples, just 100% correlations between chosen tuples of concepts and classes.
+Through SC regularization, the learned classification weights are less reliant on the revealed SCs.
+The improvement in worst group accuracy shows the new representations are more robust and better aligned with the classification task, underling that WASP identifies concepts that are truly spuriously correlated with the classes.
+For a better comparison, we replicate the SC identification process of Lg (Zhao et al., 2024) in this scenario.
+Table 4: Qualitative SCs examples, extracted on Waterbirds, CelebA and CivilComments datasets.
+See in red concepts that are off-topic, person names, or too related to the semantic content of the class, and in blue new concepts, that were not identified before.
+WASP, w.r.t.
+others, focus on learned SCs, discovering many new spuriously correlated concepts (and expressions, marked with …).
+Waterbirds CelebA CivilComments
+landbird waterbird blonde hair non-blonde hair offensive non-offensive
+B2T forest, woods, tree, branch ocean, beach, surfer, boat, dock, water, lake - man, male *github - -
+SpLiCE - - hairstyles, dolly, turban, actress, tennis, beard hairstyles, visor, amy, kate, fielder, cuff, rapper, cyclist - -
+Lg forest, woods, rainforest, tree branch, tree beach, lake, water, seagull, pond …blonde hair, actress, model, woman long hair man, man wearing sunglasses, young man, black hair, actor - -
+WASP (ours) forest…, bamboo…, ground, field, log, grass…, tree swimming…, water, lake, flying…, boat, lifeguard, pond - hat…, man…, actor, dark, person, large, shirt hypocrisy, troll, solly, hate allowing, work, made, talk
+Table 5: Zero-shot prompting with Foundation Model (FM).
+To account the learned SCs, we explicitly introduce them in the prompting scheme, comparing with those discovered by competition.
+Using WASP spurious correlated concepts, the zero-shot performance (worst group accuracy) significantly improves across all datasets (image and text modalities).
 Waterbirds (Acc % $\uparrow$) CelebA (Acc % $\uparrow$) CivilComm (Acc % $\uparrow$)
 Method Worst Avg. Worst Avg. Worst Avg.
-FM zero-shot 35.2 90.7 72.8 87.4 33.1 83.4
-FM w B2T [15] 48.1 86.1 72.8 88.0 - -
-FM w SpLiCE [4] - - 67.2 90.2 - -
-FM w Lg [47] 46.1 85.9 50.6 87.2 - -
-FM w ConceptDrift 50.3 86.3 73.1 85.7 53.2 71.0
+FM zero-shot 35.2 84.2 72.8 87.7 33.1 80.2
+FM w B2T 48.1 86.1 72.8 88.0 - -
+FM w SpLiCE - - 67.2 90.2 - -
+FM w Lg 46.1 85.9 50.6 87.2 - -
+FM w WASP 50.3 86.3 73.1 85.7 53.2 71.0
 Table 6: Ablation.
-Following the zero-shot bias-augmented prompting setup, we variate the cut-off threshold for considering causally-unrelated concepts as biases and also try to address the text-image modality gap.
+Following the zero-shot SC-augmented prompting setup, we variate the cut-off threshold for considering causally-unrelated concepts as spuriously correlated and also try to address the text-image modality gap.
 Waterbirds (Acc % $\uparrow$) CelebA (Acc % $\uparrow$) CivilComm (Acc % $\uparrow$)
@@
 modality gap: closed 48.7 85.9 72.7 85.3 - -
-ConceptDrift
+WASP
 dynamic threshold
 modality gap: open 50.3 86.3 73.1 85.7 53.2 71.0
-5.4 Bias-aware zero-shot prompting
-To further validate our identified biases, we follow Kim et al.
-[15] and evaluate them in the context of zero-shot classification task.
-We augment the initial, class-only, related prompt with the bias only through a minimal intervention (e.g.’a photo of a {cls} in the {bias}’ (see Appx.
+5.3 Spurious-aware zero-shot prompting
+To further validate our identified spurious correlations, we follow Kim et al.
+(2024) and evaluate them in the context of a zero-shot classification task.
+We augment the initial, class-oriented prompt with the identified concepts through a minimal intervention (e.g.
+’a photo of a {cls} in the {concept}’ (see Appx.
 A).
-For each class we create a prompt with each identified bias.
+For each class we create a prompt with each identified spurious correlation.
 When classifying an image, we take into account only the highest similarity among the prompts of a class (zero-shot with max-pooling over prompts).
-The results in Tab.
-5 show how the biases automatically selected by our method improve the worst group accuracy over the initial zero-shot baseline and other state-of-the-art bias-extracting solutions in all the tested datasets.
-This highlights the relevance of the biases automatically extracted by ConceptDrift.
-Ablations We validate several key decisions in our algorithm in Tab.
-6, using the zero-shot classification task and the biased-enhanced prompting.
-We notice that the number of chosen class-unrelated concepts is important, as taking too many of them adds noise to the prompts and lowers the performance.
+See in Tab.
+5 how the SCs revealed by our method improve the worst group accuracy over the initial zero-shot baseline and other state-of-the-art solutions, in all the tested datasets.
+This highlights the relevance of the SCs automatically extracted by WASP.
+5.4 Qualitative examples
+We present in Tab.
+4 the concepts identified as spuriously correlated with each class by WASP and competitors.
+Notice how our method discovers many new concepts (in blue) when compared with others.
+This is because our approach is fundamentally different, as it relies on the decision-making process of the model being investigated, diverging from current techniques oriented to validation set errors (B2T), or others that do data analysis over frozen concepts (SpLiCE, Lg).
+For CelebA-blonde hair, B2T and WASP do not found any SCs.
+This turn out to be an appropriate decision, since the presence of the feminine features do not incline the model towards one class or the other.
+See an exhaustive list of WASP revealed SCs (ImageNet-1k included) in Appx.
+C.
+Ablations We validate several WASP decisions in Tab.
+6, for zero-shot classification task, using prompts enhanced with SCs.
+The number of SCs per class turns out to be very important, taking too many adds noise to the prompts and lowers the performance.
 Nevertheless, dynamically choosing the threshold, as described in Sec.
-3-Step2c., proves to be a good strategy for adapting the cut-off across datasets.
-Following prior observations regarding the modalities gap between text and image embedding space [18, 46], we follow Liang et al.
-[18] and subtract half of the gap from the embeddings, then re-normalize them, ending in marginally lower performance w.r.t.
-not addressing the modality gap.
+3-Step2c., proves to be a good strategy for adapting the cut-off across classes.
+Following prior observations regarding the modalities gap between text and image embedding space (Liang et al., 2022; Zhang et al., 2023a), we subtract half of the gap from the embeddings and re-normalize them, ending in marginally lower performance w.r.t.
+not addressing the gap.
 6 Conclusions
-We introduce ConceptDrift, a method that identifies biases through a weight-space and filtration approach, that can be scaled to large datasets.
-By leveraging CLIP models with aligned image-text embedding, we can easily identify concepts that significantly influence class predictions, while filtering undesirable concepts.
+We introduce WASP, a method that automatically identifies SCs through a weight-space and filtration approach, that can be scaled to large datasets.
+By leveraging foundation models’ embedding space topologies, we identify concepts that significantly influence class predictions, while filtering undesirable concepts.
 We apply this method to multiple modalities on datasets like Waterbirds, CelebA, CivilComments, and ImageNet.
-We validate that the desired biases are useful by employing them in different methods for improved generalization.
-This includes a) creating subgroups that promote robust generalization b) a novel bias regularization that works even under $100\%$ correlation of biases and classes, where previous methods fail c) using better bias-aware zero-shot promoting.
-Using ConceptDrift we identify biases in ImageNet and generate synthetic images that are able to fool strong pre-trained methods, validating the relevance of the identified biases in larger datasets.
-This method opens the door for automatically finding biases in large settings without explicit human intervention.
-Supplementary Material
+We validate that the proposed SCs are useful by employing them in different methods for improved generalization.
+This includes a) a novel SC-based regularization that works even under $100\%$ correlation of concepts and classes, where previous methods fail b) using better SC-aware zero-shot promoting.
+Using WASP we identify SCs in ImageNet and generate synthetic images that are able to fool strong pre-trained methods, validating the relevance of the identified SCs in larger datasets.
+This method opens the door for automatically finding SCs in large settings without explicit human intervention.
 Appendix A Zero-Shot Prompts
 In Tab.
-A.1, we structured the class names used for initializing the initial weights of the linear layer, along with the prompt templates employed in the zero-shot classification experiments discussed in Sec.
-5.4.
-Table A.1: Class names and prompts used in the zero-shot classification task.
+7, we structured the class names used for initializing the initial weights of the linear layer, along with the prompt templates employed in the zero-shot classification experiments discussed in Sec.
+5.3.
+Table 7: Class names and prompts used in the zero-shot classification task.
 Waterbirds CelebA CivilComments
@@
 zero-shot prompt a photo of a {cls} a photo of a person with {cls} {cls}
-bias-enhanced prompt a photo of a {cls} in the {bias} a photo of a {bias} with {cls} a/an {cls} comment about {bias}
-Appendix B Biases and top class-neutral concepts
-We present the exhaustive list of biases for the Waterbirds (Tab.
-B.1 & B.2), CelebA (Tab.
-B.3 & B.4) and CivilComments (Tab.
-B.5 & B.6) datasets.
+SC-enhanced prompt a photo of a {cls} in the {SC} a photo of a {SC} with {cls} a/an {cls} comment about {SC}
+Appendix B Results of state-of-the-art models
+We provide results using the same data and experimental setup used for Tables 1 and 2, for an exhaustive list of ImageNet classifiers, in Tables 8, 9 and 10.
+Pre-trained models together with their respective weight sets are employed from the torchvision package.
+Table 8: Accuracy of various convolutional and transformer-based models trained on ImageNet-1k, on the data generated for Tab.
+1.
+As with Fig.
+1 and Fig.
+4, we note that the performance of these models is significantly affected, even though the correct class is clearly illustrated right in front and center while, and the predicted class is absent from the generated images.
+Prompt employed (correct class highlighted in bold, SC in red)
+Model - Weights a photo of a peafowl firemen and a peafowl a photo of a Bernese Mountain Dog shrimp and pasta near a Bernese Mountain Dog
+alexnet - V1 (Krizhevsky, 2014) 100.0 4.6 96.2 23.3
+convnext_tiny - V1 (Liu et al., 2022) 100.0 77.2 94.2 69.8
+convnext_small - V1 (Liu et al., 2022) 100.0 92.9 96.3 78.3
+convnext_base - V1 (Liu et al., 2022) 100.0 83.5 99.3 78.0
+convnext_large - V1 (Liu et al., 2022) 100.0 88.2 99.6 81.9
+densenet121 - V1 (Huang et al., 2016) 100.0 52.3 93.1 74.8
+densenet161 - V1 (Huang et al., 2016) 100.0 52.8 84.0 51.6
+densenet201 - V1 (Huang et al., 2016) 100.0 49.7 86.1 80.8
+efficientnet_b0 - V1 (Tan & Le, 2019) 100.0 64.5 99.2 92.9
+efficientnet_b1 - V1 (Tan & Le, 2019) 100.0 42.6 88.1 67.1
+efficientnet_b1 - V2 (Tan & Le, 2019) 100.0 84.2 99.9 69.5
+efficientnet_b2 - V1 (Tan & Le, 2019) 100.0 61.7 99.6 82.4
+efficientnet_b3 - V1 (Tan & Le, 2019) 100.0 89.1 99.1 92.8
+efficientnet_b4 - V1 (Tan & Le, 2019) 100.0 94.4 99.9 72.7
+efficientnet_b5 - V1 (Tan & Le, 2019) 100.0 82.9 99.4 86.3
+efficientnet_b6 - V1 (Tan & Le, 2019) 100.0 91.2 100.0 95.6
+efficientnet_b7 - V1 (Tan & Le, 2019) 100.0 88.3 99.9 90.3
+efficientnet_v2_s - V1 (Tan & Le, 2021) 100.0 98.3 99.2 88.1
+efficientnet_v2_m - V1 (Tan & Le, 2021) 100.0 95.1 99.8 94.3
+efficientnet_v2_l - V1 (Tan & Le, 2021) 100.0 96.0 99.1 83.6
+googlenet - V1 (Szegedy et al., 2015) 100.0 45.4 95.2 57.0
+inception_v3 - V1 (Szegedy et al., 2016) 100.0 82.2 98.9 80.8
+maxvit_t - V1 (Tu et al., 2022) 100.0 91.7 99.7 85.3
+mnasnet0_5 - V1 (Tan et al., 2019) 100.0 23.3 96.7 60.8
+mnasnet0_75 - V1 (Tan et al., 2019) 100.0 36.9 98.4 71.1
+mnasnet1_0 - V1 (Tan et al., 2019) 100.0 32.0 89.4 74.9
+mnasnet1_3 - V1 (Tan et al., 2019) 100.0 63.9 91.7 75.6
+mobilenet_v2 - V1 (Sandler et al., 2018) 100.0 26.2 84.0 45.0
+mobilenet_v2 - V2 (Sandler et al., 2018) 100.0 54.1 98.6 62.8
+mobilenet_v3_small - V1 (Howard et al., 2019) 100.0 16.6 91.0 30.3
+mobilenet_v3_large - V1 (Howard et al., 2019) 100.0 25.4 94.2 23.1
+mobilenet_v3_large - V2 (Howard et al., 2019) 100.0 56.0 96.6 73.7
+Table 9: Accuracy of various convolutional and transformer-based models trained on ImageNet-1k, on the data generated for Tab.
+1.
+As with Fig.
+1 and Fig.
+4, we note that the performance of these models is significantly affected, even though the correct class is clearly illustrated right in front and center while, and the predicted class is absent from the generated images.
+Prompt employed (correct class highlighted in bold, SC in red)
+Model - Weights a photo of a peafowl firemen and a peafowl a photo of a Bernese Mountain Dog shrimp and pasta near a Bernese Mountain Dog
+regnet_y_400mf - V1 (Radosavovic et al., 2020) 100.0 35.6 74.7 34.7
+regnet_y_400mf - V2 (Radosavovic et al., 2020) 100.0 72.1 97.7 87.9
+regnet_y_800mf - V1 (Radosavovic et al., 2020) 100.0 22.7 94.8 54.4
+regnet_y_800mf - V2 (Radosavovic et al., 2020) 100.0 81.7 98.9 88.0
+regnet_y_1_6gf - V1 (Radosavovic et al., 2020) 100.0 47.0 96.4 71.7
+regnet_y_1_6gf - V2 (Radosavovic et al., 2020) 100.0 88.2 96.4 74.1
+regnet_y_3_2gf - V1 (Radosavovic et al., 2020) 100.0 33.5 99.4 90.9
+regnet_y_3_2gf - V2 (Radosavovic et al., 2020) 100.0 94.6 98.4 75.4
+regnet_y_8gf - V1 (Radosavovic et al., 2020) 100.0 58.3 71.3 54.8
+regnet_y_8gf - V2 (Radosavovic et al., 2020) 100.0 98.2 96.8 67.5
+regnet_y_16gf - V1 (Radosavovic et al., 2020) 100.0 86.7 97.8 49.8
+regnet_y_16gf - V2 (Radosavovic et al., 2020) 100.0 98.7 91.8 83.5
+regnet_y_16gf - SWAG_E2E_V1 (Radosavovic et al., 2020) 100.0 99.6 97.7 78.2
+regnet_y_16gf - SWAG_LINEAR_V1 (Radosavovic et al., 2020) 100.0 78.3 100.0 90.7
+regnet_y_32gf - V1 (Radosavovic et al., 2020) 100.0 84.5 99.1 82.3
+regnet_y_32gf - V2 (Radosavovic et al., 2020) 100.0 98.0 99.5 79.6
+regnet_y_32gf - SWAG_E2E_V1 (Radosavovic et al., 2020) 100.0 99.6 93.5 71.5
+regnet_y_32gf - SWAG_LINEAR_V1 (Radosavovic et al., 2020) 100.0 97.2 100.0 85.6
+regnet_y_128gf - SWAG_E2E_V1 (Radosavovic et al., 2020) 100.0 99.6 54.4 67.8
+regnet_y_128gf - SWAG_LINEAR_V1 (Radosavovic et al., 2020) 100.0 90.8 99.9 96.9
+regnet_x_400mf - V1 (Radosavovic et al., 2020) 100.0 37.4 82.3 29.0
+regnet_x_400mf - V2 (Radosavovic et al., 2020) 100.0 50.7 99.5 81.7
+regnet_x_800mf - V1 (Radosavovic et al., 2020) 100.0 25.4 77.0 52.6
+regnet_x_800mf - V2 (Radosavovic et al., 2020) 100.0 75.5 97.1 71.4
+regnet_x_1_6gf - V1 (Radosavovic et al., 2020) 100.0 38.2 76.9 66.7
+regnet_x_1_6gf - V2 (Radosavovic et al., 2020) 100.0 82.2 99.2 88.1
+regnet_x_3_2gf - V1 (Radosavovic et al., 2020) 100.0 45.1 62.1 69.4
+regnet_x_3_2gf - V2 (Radosavovic et al., 2020) 100.0 83.4 99.6 88.3
+regnet_x_8gf - V1 (Radosavovic et al., 2020) 100.0 41.8 98.8 89.0
+regnet_x_8gf - V2 (Radosavovic et al., 2020) 100.0 93.5 99.2 81.6
+regnet_x_16gf - V1 (Radosavovic et al., 2020) 100.0 48.7 86.6 54.5
+regnet_x_16gf - V2 (Radosavovic et al., 2020) 100.0 93.4 97.9 88.1
+regnet_x_32gf - V1 (Radosavovic et al., 2020) 100.0 66.1 85.9 46.0
+regnet_x_32gf - V2 (Radosavovic et al., 2020) 100.0 97.4 99.6 83.7
+resnet18 - V1 (He et al., 2016) 100.0 36.8 84.3 41.9
+resnet34 - V1 (He et al., 2016) 100.0 32.9 54.1 26.1
+resnet50 - V1 (He et al., 2016) 100.0 30.1 73.9 54.5
+resnet50 - V2 (He et al., 2016) 100.0 80.1 99.7 88.4
+resnet101 - V1 (He et al., 2016) 100.0 60.4 92.3 84.4
+resnet101 - V2 (He et al., 2016) 100.0 93.4 98.8 87.4
+resnet152 - V1 (He et al., 2016) 100.0 66.1 98.4 78.2
+resnet152 - V2 (He et al., 2016) 100.0 93.1 98.5 92.5
+resnext50_32x4d - V1 (Xie et al., 2017) 100.0 45.5 92.6 74.8
+resnext50_32x4d - V2 (Xie et al., 2017) 100.0 80.3 98.5 88.0
+resnext101_32x8d - V1 (Xie et al., 2017) 100.0 66.6 84.7 61.2
+resnext101_32x8d - V2 (Xie et al., 2017) 100.0 90.3 99.4 85.2
+resnext101_64x4d - V1 (Xie et al., 2017) 100.0 77.9 97.7 74.2
+Table 10: Accuracy of various convolutional and transformer-based models trained on ImageNet-1k, on the data generated for Tab.
+1.
+As with Fig.
+1 and Fig.
+4, we note that the performance of these models is significantly affected, even though the correct class is clearly illustrated right in front and center while, and the predicted class is absent from the generated images.
+Prompt employed (correct class highlighted in bold, SC in red)
+Model - Weights a photo of a peafowl firemen and a peafowl a photo of a Bernese Mountain Dog shrimp and pasta near a Bernese Mountain Dog
+shufflenet_v2_x0_5 - V1 (Ma et al., 2018) 100.0 23.8 36.9 21.0
+shufflenet_v2_x1_0 - V1 (Ma et al., 2018) 99.8 30.4 72.1 55.9
+shufflenet_v2_x1_5 - V1 (Ma et al., 2018) 100.0 41.2 97.9 52.5
+shufflenet_v2_x2_0 - V1 (Ma et al., 2018) 100.0 61.4 99.3 64.5
+squeezenet1_0 - V1 (Iandola et al., 2016) 100.0 11.4 95.9 28.7
+squeezenet1_1 - V1 (Iandola et al., 2016) 100.0 13.8 91.2 46.1
+swin_t - V1 (Liu et al., 2021b) 100.0 72.7 96.9 81.6
+swin_s - V1 (Liu et al., 2021b) 100.0 74.3 99.3 81.4
+swin_b - V1 (Liu et al., 2021b) 100.0 81.5 95.2 72.6
+swin_v2_t - V1 (Liu et al., 2021b) 100.0 76.2 88.7 73.6
+swin_v2_s - V1 (Liu et al., 2021b) 100.0 85.7 90.7 74.4
+swin_v2_b - V1 (Liu et al., 2021b) 100.0 73.0 96.0 85.2
+vgg11 - V1 (Simonyan & Zisserman, 2015) 100.0 9.9 96.6 60.7
+vgg11_bn - V1 (Simonyan & Zisserman, 2015) 100.0 15.9 86.7 54.8
+vgg13 - V1 (Simonyan & Zisserman, 2015) 100.0 5.9 97.7 68.1
+vgg13_bn - V1 (Simonyan & Zisserman, 2015) 100.0 15.5 87.0 13.0
+vgg16 - V1 (Simonyan & Zisserman, 2015) 100.0 12.2 93.7 66.0
+vgg16_bn - V1 (Simonyan & Zisserman, 2015) 100.0 22.6 96.1 70.7
+vgg19 - V1 (Simonyan & Zisserman, 2015) 100.0 26.6 98.5 40.8
+vgg19_bn - V1 (Simonyan & Zisserman, 2015) 100.0 35.9 83.1 46.2
+vit_b_16 - V1 (Dosovitskiy et al., 2021) 100.0 85.4 97.1 75.2
+vit_b_16 - SWAG_E2E_V1 (Dosovitskiy et al., 2021) 100.0 88.9 95.8 59.1
+vit_b_16 - SWAG_LINEAR_V1 (Dosovitskiy et al., 2021) 100.0 79.2 99.9 84.6
+vit_b_32 - V1 (Dosovitskiy et al., 2021) 100.0 56.1 96.0 86.0
+vit_l_16 - V1 (Dosovitskiy et al., 2021) 100.0 55.9 95.3 76.0
+vit_l_16 - SWAG_E2E_V1 (Dosovitskiy et al., 2021) 100.0 92.1 100.0 93.3
+vit_l_16 - SWAG_LINEAR_V1 (Dosovitskiy et al., 2021) 100.0 97.4 100.0 72.5
+vit_l_32 - V1 (Dosovitskiy et al., 2021) 100.0 54.0 96.9 82.9
+vit_h_14 - SWAG_E2E_V1 (Dosovitskiy et al., 2021) 100.0 99.4 98.5 86.4
+vit_h_14 - SWAG_LINEAR_V1 (Dosovitskiy et al., 2021) 100.0 99.7 100.0 96.1
+wide_resnet50_2 - V1 (Zagoruyko & Komodakis, 2016) 100.0 60.6 95.7 63.9
+wide_resnet50_2 - V2 (Zagoruyko & Komodakis, 2016) 100.0 83.9 99.5 85.4
+wide_resnet101_2 - V1 (Zagoruyko & Komodakis, 2016) 100.0 69.3 84.1 72.8
+wide_resnet101_2 - V2 (Zagoruyko & Komodakis, 2016) 100.0 91.0 98.8 84.9
+Appendix C SCs and top class-neutral concepts
+We present the exhaustive list of SCs found for the Waterbirds (Tab.
+12 & 12), CelebA (Tab.
+13) and CivilComments (Tab.
+15 & 15) datasets.
 We also present top class-neutral concepts for ImageNet classes.
@@
 Accounting for the size of the dataset, we will publish the available data on our repository upon acceptance, and we will restrict the presentation within the context of the current format to a few classes for illustrative purposes in Tab.
-B.7 -B.15.
-Table B.1: Top Waterbirds class-neutral concepts for ”landbird”.
+20 -19.
+Table 11: Top Waterbirds class-neutral concepts for ”landbird”.
 Landbird Score
@@
 hill 0.03341371
-Table B.2: Top Waterbirds class-neutral concepts for ”waterbird”.
+Table 12: Top Waterbirds class-neutral concepts for ”waterbird”.
 Waterbird Score
@@
 pond 0.053261578
-Table B.3: Top CelebA class-neutral concepts for ”non-blonde”.
+Table 13: Top CelebA class-neutral concepts for ”non-blonde”.
 Non-Blonde Score
@@
 man face 0.12130207
-Table B.4: Top CelebA class-neutral concepts for ”blonde”.
-Blonde Score
-yellow -0.021363914
-white -0.04120779
-golden -0.041870475
-woman with blowing -0.042802572
-woman with laughing -0.046326935
-girl with blue eyes -0.050636113
-woman with playing -0.0510059
-girl next door -0.05101496
-woman with green -0.05205536
-sweater -0.054968417
-gold -0.057154775
-scarf -0.0585289
-pale -0.05990547
-woman with blue eyes -0.05999875
-tank top -0.060728252
-Table B.5: Top CivilComments class-neutral concepts for ”non-offensive”.
+Table 14: Top CivilComments class-neutral concepts for ”non-offensive”.
 Non-offensive Score
@@
 attempt 0.056797385
-Table B.6: Top CivilComments class-neutral concepts for ”offensive”.
+Table 15: Top CivilComments class-neutral concepts for ”offensive”.
 ”Offensive” – Top Concepts Score
@@
 kid you have the absolute 0.001619577
-die on the golf -0.001492977
-Muslim Canadians Party -0.002062559
-serve them at the lunch -0.00358063
-culture is no worse -0.005233765
-banned -0.006861687
-pretend they are all goodness -0.007072449
-pop that particular rainbow -0.008466244
-member becomes so sick -0.008648455
-men and she wanted -0.008733362
-Canadian Muslims -0.009060442
-destroy -0.010027111
-six Baltic -0.01069364
-customers when it not acceptable -0.011366904
-Table B.7: Top ImageNet class-neutral concepts for ”American Lobster”.
-”American Lobster” – Top Concepts Score
-pasta with shrimp 0.3148532
-adirondack sitting 0.31178916
-large shrimp 0.3085378
-shrimp and pasta 0.30821544
-shrimp cooking 0.3032636
-large cast cooking 0.30282205
-pasta with shrimp and cheese 0.2984723
-adirondack 0.29169592
-and mussels 0.2904386
-legs and other seafood 0.28870153
-close up of a shrimp 0.2871693
-of pasta with shrimp 0.2823377
-seafood 0.2808096
-mussels 0.28018713
-cast cooking 0.27951774
-shrimp and cheese 0.27841187
-shrimp 0.2726224
-Table B.8: Top ImageNet class-neutral concepts for ”Bald Eagle”.
-”Bald Eagle” – Top Concepts Score
-osprey flying 0.2888234
-osprey 0.27377927
-row of american flags 0.26513425
-group of american flags flying 0.25681192
-american flag and american flag 0.23945728
-emu standing 0.23752406
-yellowstone national 0.23740456
-Table B.9: Top ImageNet class-neutral concepts for ”Ballpoint Pen”.
-”Ballpoint Pen” – Top Concepts Score
-wearing a pilot 0.36252645
-markers 0.36050144
-sharpie 0.34563553
-stylus 0.34399948
-pair of eyeglasses 0.32426757
-notepad 0.3228797
-calligraphy 0.3222967
-paint and markers 0.32198787
-close up of a needle 0.32131955
-on a straw 0.3209229
-and markers 0.32090995
-eyeglasses 0.31017447
-dots 0.30764964
-crayons 0.30598855
-Table B.10: Top ImageNet class-neutral concepts for ”Crossword”.
+Table 16: Top ImageNet class-neutral concepts for ”Crossword”.
 ”Crossword” – Top Concepts Score
@@
 square of crochet 0.22324148
-Table B.11: Top ImageNet class-neutral concepts for ”Coffeemaker”.
+Table 17: Top ImageNet class-neutral concepts for ”Guacamole”.
+”Guacamole” – Top Concepts Score
+tomatoes and avocado 0.33948907
+avocado 0.3125001
+nachos 0.30761188
+ham and parsley 0.2960047
+with avocado 0.2952027
+colorful mexican 0.28313732
+bacon and parsley 0.27773544
+of avocado 0.2754345
+side of salsa 0.27470407
+peridot 0.2734925
+salsa 0.27124816
+nachos with 0.27092364
+lime body 0.27088284
+cheese and parsley 0.26951405
+pile of limes 0.26944226
+peas and bacon 0.2671204
+limes and limes 0.26682717
+lime cut 0.2662533
+nachos with and 0.26580203
+pasta with peas 0.2626204
+tacos 0.2619322
+pasta with ham and parsley 0.26183394
+tomatoes and cilantro 0.26161948
+Table 18: Top ImageNet class-neutral concepts for ”Bald Eagle”.
+”Bald Eagle” – Top Concepts Score
+osprey flying 0.2888234
+osprey 0.27377927
+row of american flags 0.26513425
+group of american flags flying 0.25681192
+american flag and american flag 0.23945728
+emu standing 0.23752406
+yellowstone national 0.23740456
+Table 19: Top ImageNet class-neutral concepts for ”Ballpoint Pen”.
+”Ballpoint Pen” – Top Concepts Score
+wearing a pilot 0.36252645
+markers 0.36050144
+sharpie 0.34563553
+stylus 0.34399948
+pair of eyeglasses 0.32426757
+notepad 0.3228797
+calligraphy 0.3222967
+paint and markers 0.32198787
+close up of a needle 0.32131955
+on a straw 0.3209229
+and markers 0.32090995
+eyeglasses 0.31017447
+dots 0.30764964
+crayons 0.30598855
+Table 20: Top ImageNet class-neutral concepts for ”American Lobster”.
+”American Lobster” – Top Concepts Score
+pasta with shrimp 0.3148532
+adirondack sitting 0.31178916
+large shrimp 0.3085378
+shrimp and pasta 0.30821544
+shrimp cooking 0.3032636
+large cast cooking 0.30282205
+pasta with shrimp and cheese 0.2984723
+adirondack 0.29169592
+and mussels 0.2904386
+legs and other seafood 0.28870153
+close up of a shrimp 0.2871693
+of pasta with shrimp 0.2823377
+seafood 0.2808096
+mussels 0.28018713
+cast cooking 0.27951774
+shrimp and cheese 0.27841187
+shrimp 0.2726224
+Table 21: Top ImageNet class-neutral concepts for ”Coffeemaker”.
 ”Coffeemaker” – Top Concepts Score
@@
 vases sitting 0.3643943
-Table B.12: Top ImageNet class-neutral concepts for ”Doormat”.
+Table 22: Top ImageNet class-neutral concepts for ”Doormat”.
 ”Doormat” – Top Concepts Score
@@
 standing in a doorway 0.3126963
-Table B.13: Top ImageNet class-neutral concepts for ”Eraser”.
+Table 23: Top ImageNet class-neutral concepts for ”Eraser”.
 ”Eraser” – Top Concepts Score
@@
 band aid flexible 0.32946587
-Table B.14: Top ImageNet class-neutral concepts for ”Fire Truck”.
+Table 24: Top ImageNet class-neutral concepts for ”Fire Truck”.
 ”Fire Truck” – Top Concepts Score
@@
 flames painted 0.2540929
-Table B.15: Top ImageNet class-neutral concepts for ”Guacamole”.
-”Guacamole” – Top Concepts Score
-tomatoes and avocado 0.33948907
-avocado 0.3125001
-nachos 0.30761188
-ham and parsley 0.2960047
-with avocado 0.2952027
-colorful mexican 0.28313732
-bacon and parsley 0.27773544
-of avocado 0.2754345
-side of salsa 0.27470407
-peridot 0.2734925
-salsa 0.27124816
-nachos with 0.27092364
-lime body 0.27088284
-cheese and parsley 0.26951405
-pile of limes 0.26944226
-peas and bacon 0.2671204
-limes and limes 0.26682717
-lime cut 0.2662533
-nachos with and 0.26580203
-pasta with peas 0.2626204
-tacos 0.2619322
-pasta with ham and parsley 0.26183394
-tomatoes and cilantro 0.26161948
-Appendix C Algorithm
-We present the formal implementation of our proposed ConceptDrift approach in Alg.
+Appendix D Algorithm
+We present the formal implementation of our proposed WASP approach in Alg.
 1.
@@
 The filtering can be performed by means of employing WordNet associations, Large Language Models, or both.
-With the embeddings of the class-neutral concepts at hand, we proceed to determine the biases for each class.
-At line 7, we initialize our set of biases with an empty set, and we compute the similarities between each selected concept and each class at line 8.
+With the embeddings of the class-neutral concepts at hand, we proceed to determine the SCs for each class.
+At line 7, we initialize our set of spuriously correlated concepts with an empty set, and we compute the similarities between each selected concept and each class at line 8.
 We further proceed to iterate through each class (line 9).
@@
 3.
-Finally, we select the concepts with ranking scores above the computed threshold and store them as biases for class $k$, at lines 15 and 16.
-Algorithm 1 ConceptDrift
+Finally, we select the concepts with ranking scores above the computed threshold and store them as spuriously correlated concepts for class $k$, at lines 15 and 16.
+Algorithm 1 WASP
 Input:
@@
 r - window size for dynamic thresholding smoothing
-Output: Identified positively correlated biases $\mathcal{B}$.
+Output: Identified positive SCs $\mathcal{B}$.
 1:   // Step 1. Model Training
@@
 3:  $\mathbf{W}\leftarrow\mbox{ERM}\left(\mathbf{W^{0}},\mbox{M},(\mathcal{X},\mathcal{Y}),(\mathcal{X}_{val},\mathcal{Y}_{val})\right)$
-4:    // Step 2. Bias identification
+4:   // Step 2. SC identification
 5:  $C\leftarrow$Filter$(C_{all})$ // 2a. Filter concepts: LLM/WordNet
@@
 15:     $b_{k}=\mathcal{C}\left[s_{k}\geq th_{k}\right]$
-16:      $\mathcal{B}=\mathcal{B}\cup(k,b_{k})$ // Positively correlated biases
-17:   end for
-18:   return  $\mathcal{B}$
-Appendix D Loss to bias similarity correlation
-To validate that the environments discovered through the partitioning approach from B2T [15], using the biases uncovered with our method, are indeed effective in reducing the model’s reliance on shortcuts, we analyzed the correlation between the similarity of each training image to biases and its training loss value under both ERM and GroupDRO training, using our identified groups.
-The results show a reduction in correlation scores across all biases, demonstrating that the revealed groups are relevant to the dataset’s underlying distribution, and can be effectively utilized with specific algorithms to mitigate the model’s dependence on spurious correlations.
+16:     $\mathcal{B}=\mathcal{B}\cup(k,b_{k})$ // Positive SCs
+17:  end forReturn $\mathcal{B}$
+Appendix E Loss correlation with presence of spuriously correlated concepts
+In this experiment, we look at the correlations between the loss values and the concept-to-sample similarities.
+We compare basic ERM with GroupDRO, applied on groups, that are obtained based on our revealed SCs (and further grouped using the B2T (Kim et al., 2024) partitioning strategy).
+See in Fig.
+5 how for GroupDRO, the loss-to-similarity correlation significantly decrease, revealing that the model is less prone to make mistakes on the samples containing SCs.
+The results show a reduction in correlation scores across all SCs, demonstrating that the revealed groups are relevant to the dataset’s underlying distribution, and can be effectively utilized with specific algorithms to mitigate the model’s dependence on spurious correlations.
 Fig.
-D.1 shows the Pearson correlation scores after an epoch of training on Waterbirds, on a subset of all biases.
-Figure D.1: Correlation(sample_loss, sample_to_bias similarity) under ERM/GDRO after one epoch of training on Waterbirds.
+5 shows the Pearson correlation scores after an epoch of training on Waterbirds, on a subset of all SCs.
+Figure 5: Correlation(sample_loss, sample_to_bias similarity) under ERM/GDRO after one epoch of training on Waterbirds.
 Loss correlation w/ biases, ERM vs GroupDRO using groups created with the B2T partitioning method.